### 11 Mar 2024 v2

# Integral cohomology rings of weighted Grassmann orbifolds and rigidity properties

Abstract.
In this paper, we introduce ‘Plücker weight vector’ and establish the definition of a weighted Grassmann orbifold $\mbox{Gr}_{\mathbf{b}}(k,n)$, corresponding to a Plücker weight vector ‘$\mathbf{b}$’. We achieve an explicit classification of weighted Grassmann orbifolds up to certain homeomorphism in terms of the Plücker weight vectors. We study the integral cohomology of $\mbox{Gr}_{\bf b}(k,n)$ and provide some sufficient conditions such that the integral cohomology of $\mbox{Gr}_{\bf b}(k,n)$ has no torsion. We describe the integral equivariant cohomology ring of divisive weighted Grassmann orbifolds and compute all the equivariant structure constants with integer coefficients. Eminently, we compute the integral cohomology rings of divisive weighted Grassmann orbifolds explicitly.
Key words and phrases:
Plücker weight vector, weighted Grassmann orbifold, $q$-CW complex, Plücker permutation, lens space, integral cohomology, integral equivariant cohomology, equivariant Schubert basis, structure constant
2020 Mathematics Subject Classification:
14M15, 57R18, 55N10, 55N91
1. Introduction
Grassmann manifolds are the central objects of study in algebraic topology, algebraic geometry, and enumerative geometry, see [12, 14, 15]. Consider the $n$-dimensional complex vector space $\mathbb{C}^{n}$ and a positive integer $k$ satisfying $1\leq k<n$. The space comprising all $k$-dimensional subspaces in the complex $n$-dimensional space $\mathbb{C}^{n}$ possesses the structure of a smooth manifold. This manifold is known as a complex Grassmann manifold (in short, Grassmann manifold in this paper) and is denoted by $\mbox{Gr}(k,n)$. Grassmann manifold $\mbox{Gr}(k,n)$ can be equivariantly embedded into the complex projective space $\mathbb{C}P^{m}$, where $m+1={n\choose k}$. This embedding is known as the Plücker embedding, presenting $\mbox{Gr}(k,n)$ as a projective algebraic variety. The image of the Grassmann manifold in the projective space is described by a set of homogeneous polynomial equations known as Plücker relations. The non-zero points in $\mathbb{C}^{m+1}$ that satisfy Plücker relations are referred to as Plücker coordinates.
In [13], Kawasaki introduced a ${\bf{b}}$-action of $\mathbb{C}^{*}$ on $\mathbb{C}^{m+1}\setminus\{\mathbf{0}\}$ for a weight vector ${\bf{b}}=(b_{0},\dots,b_{m})\in(\mathbb{Z}_{\geq 1})^{m+1}$ as follows:
(1.1) $t(z_{0},z_{1},z_{2},\dots,z_{m})=(t^{b_{0}}z_{0},t^{b_{1}}z_{1},t^{b_{2}}z_{2},\dots,t^{b_{m}}z_{m}).$
The weighted projective space $\mathbb{W}P(b_{0},b_{1},\dots,b_{m})$ is defined as the orbit space of the ${\bf{b}}$-action of $\mathbb{C}^{*}$ on $\mathbb{C}^{m+1}\setminus\{\mathbf{0}\}$. Notably, when ${\bf{b}}$ is chosen as $(1,1,\dots,1)$, it becomes the complex projective space $\mathbb{C}P^{m}$. In the article [13], Kawasaki proved that the integral cohomologies of weighted projective spaces have no torsion and are concentrated in even degrees. Moreover, he computed the integral cohomology ring of weighted projective spaces explicitly. In [10], Harada, Holm, Ray, and Williams computed the integral generalized equivariant cohomology ring of a divisive weighted projective space. Several cohomological properties of weighted projective space can be found in [3], and [6].
Corti and Reid introduced weighted Grassmannian in [8] as the weighted projective analog of the Grassmann manifold. For any two positive integers $k<n$, Abe and Matsumura defined weighted Grassmannian $\mbox{WGr}(k,n)$ in [1] corresponding to a weight vector $W=(w_{0},w_{1},\dots,w_{n})\in(\mathbb{Z}_{\geq 0})^{n}$ and $a\in\mathbb{Z}_{\geq 1}$. They studied the equivariant cohomology ring of $\mbox{WGr}(k,n)$ with rational coefficients. More generally, the equivariant cohomology ring of weighted Flag varieties with rational coefficients was studied in [4]. Hilbert series of some weighted flag varieties was studied in [17, 18]. The author and Sarkar [7] gave another topological definition of $\mbox{WGr}(k,n)$ and called them weighted Grassmann orbifolds. It has an orbifold structure and admits a $q$-CW complex structure similar to the Schubert cell decomposition of the Grassmann manifold. For basic properties of orbifolds, readers are referred to [2].
In this article, we delve into the concept of Plücker coordinates, which are uniquely defined by a set of homogeneous polynomial equations referred to as Pl”ucker relations, as illustrated in (2.5). This exploration leads us to introduce Plücker weight vector ${\bf{b}}=(b_{0},b_{1},\dots,b_{m})\in(\mathbb{Z}_{\geq 0})^{m+1}$, where $m+1={n\choose k}$ for two positive integers $k$ and $n$ such that $k<n$. Then corresponding to a Plücker weight vector ${\bf{b}}$, we define a topological space $\mbox{Gr}_{\bf{b}}(k,n)$. We derive that the definition of $\mbox{Gr}_{\bf b}(k,n)$ generalizes the definition of weighted Grassmannian in [1] as well as weighted Grassmann orbifold in [7]. We show that $\mbox{Gr}_{\bf b}(k,n)$ has an orbifold structure and call this a weighted Grassmann orbifold. This broadens the class of weighted Grassmann orbifolds in terms of Plücker weight vectors. We introduce Plücker permutation, which produces several invariant $q$-CW complex structures on $\mbox{Gr}_{\bf b}(k,n)$. Moreover, we characterize weighted Grassmann orbifolds up to certain homeomorphism in terms of Plücker weight vectors and Plücker permutations. We discuss the integral cohomology of weighted Grassmann orbifolds. Inspired by the work of [13], one can ask the following.
Question 1.1.
Is the integral cohomology of a weighted Grassmann orbifold concentrated in even degrees without torsion? If this happens, can we calculate the integral cohomology rings of weighted Grassmann orbifolds?
The author and Sarkar [7] gave an affirmative answer to the first part of this question for divisive $\mbox{WGr}(k,n)$ and computed the generalized $T^{n}$-equivariant cohomology ring of divisive $\mbox{WGr}(k,n)$, where $T^{n}$ is the standard compact abelian torus of dimension $n$. Moreover, they found an equivariant Schubert basis for the equivariant cohomology ring $H_{T^{n}}^{*}(\mbox{WGr}(k,n);\mathbb{Z})$ as $H^{*}_{T^{n}}(\{pt\};\mathbb{Z})$ module. In this article, we explicitly calculate the equivariant structure constants with respect to the equivariant Schubert basis in $H_{T^{n}}^{*}(\mbox{Gr}_{\bf b}(k,n);\mathbb{Z})$ with integer coefficients. This paves the way for computing the cup product for the cohomology ring of a divisive weighted Grassmann orbifold with integer coefficients. The paper is organized as follows.
In Section 2, we explore the reltionship between the Plücker relations and Plücker coordinates. We denote the set of all Plücker coordinates by $Pl(k,n)$ as a subset of $\mathbb{C}^{m+1}\setminus\{\mathbf{0}\}$, where $m+1={n\choose k}$. The Plücker relations inspired us to introduce Plücker weight vector ${\bf{b}}=(b_{0},b_{1},\dots,b_{m})\in(\mathbb{Z}_{\geq 1})^{m+1}$ in the manner that $Pl(k,n)$ is invariant subset of $\mathbb{C}^{m+1}\setminus\{\mathbf{0}\}$ with respect to the ${\bf{b}}$-action of $\mathbb{C}^{*}$ defined in (1.1), where $m+1={n\choose k}$. Then we define a topological space $\mbox{Gr}_{\bf{b}}(k,n)$, as the orbit space of $Pl(k,n)$ by the ${\bf b}$-action of $\mathbb{C}^{*}$ for a Plücker weight vector ${\bf b}=(b_{0},b_{1},\dots,b_{m})$, see Definition 2.6. We prove that this class of $\mbox{Gr}_{\bf b}(k,n)$ broadens the class of weighted Grasmann orbifold discussed in [1] and [7]; see Proposition 2.9. We discuss the orbifold and $q$-CW complex structure of $\mbox{Gr}_{\bf{b}}(k,n)$ and call it a weighted Grassmann orbifold. The $q$-CW complex structure of $\mbox{Gr}_{\bf{b}}(k,n)$ induces a building sequence
$\{pt\}=X^{0}\subset X^{1}\subset X^{2}\subset\dots\subset X^{m}=\mbox{Gr}_{\bf b}(k,n)$
of $\mbox{Gr}_{\bf{b}}(k,n)$, see (2.12). This induces a cofibration
$L^{\prime}(b_{i};{\bf{b}}^{(i)})\hookrightarrow X^{i-1}\to X^{i}$
for the lens complex $L^{\prime}(b_{i};{\bf{b}}^{(i)})$, see Proposition 2.12.
In Section 3, we introduce Plücker permutations $\sigma$ on $\{0,1,\dots,m\}$ those permute the Plücker coordinates. Then we show that all Plücker weight vectors are stable under Plücker permutations. We prove that two weighted Grassmann orbifolds are weakly equivariantly homeomorphic if their Plücker weight vectors differ by a Plücker permutation, see Theorem 3.4. This provides several different $q$-CW complex structures of a weighted Grassmann orbifold. Conversely, we prove the following.
Theorem A (Theorem 3.11).
If there exists a homeomorphism $f$ between two weighted Grassmann orbifolds ${\rm{Gr}}_{\bf b}(k,n)$ and ${\rm{Gr}}_{\bf c}(k,n)$ such that the $f$ maps every coordinate element to some coordinate element then ${\bf b}$ and ${\bf c}$ are same up to a scalar multiplication and a permutation.
The coordinate elements are the set of all fixed points for the several group actions on $\mbox{Gr}_{\bf b}(k,n)$. Thus the above Theorem is classifies the weighted Grassmann orbifolds up to weakly equivariant homeomorphism, see Corollary 3.12.
In Section 4, for a prime $p$, we deduce the integral cohomology of any weighted Grassmann orbifold has no $p$-torsion under some hypothesis on the Plücker weighted vector, see Theorem 4.1. Moreover, we prove thae following.
Theorem B (Theorem 4.5).
$H^{i}({\rm{Gr}}_{\bf{b}}(2,4);\mathbb{Z})$ is torsion-free and concentrated in even degrees for any weighted Grassmann orbifold ${\rm{Gr}}_{\bf b}(2,4)$ if $i\neq 3$.
We study some non-trivial examples of $\mbox{Gr}_{\bf{b}}(2,4)$ and $\mbox{Gr}_{\bf{b}}(2,5)$ such that the integral cohomologies of those weighted Grassmann orbifolds have no torsion and concentrated in even degrees. We define divisive weighted Grassmann orbifolds in terms of Plücker permutations and Plücker weight vectors. Then we show that the integral cohomology of divisive Grassmann orbifolds has no torsion and is concentrated in even degrees.
In Section 5, we describe the $T^{n}$-equivariant cohomology ring of divisive $\mbox{Gr}_{\bf{b}}(k,n)$ with integer coefficients as well as a basis $\{{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{m}$ for the aforementioned ring as a $\mathbb{Z}[y_{1},\dots,y_{n}]$ module following [7]. Subsequently, we can define the equivariant structure constant ${\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{\ell}$ satisfies the following.
(1.2) ${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{j}}=\sum_{\ell=0}^{m}{\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{\ell}{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{\ell}}$
If ${\bf{b}}=(1,1,\dots,1)$, the above basis leads to a basis $\{\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{m}$ for $H_{T^{n}}^{*}(\mbox{Gr}(k,n);\mathbb{Z})$. We find the relation between $\{\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{m}$ and the basis considered by Knutson and Tao in [14] for the equivariant cohomology ring $H_{T^{n}}^{*}(\mbox{Gr}(k,n);\mathbb{Z})$. Initially we compute the equivariant structure constants $\widetilde{\mathscr{C}}_{i~{}j}^{\ell}$ with respect to the basis $\{\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{m}$, see Lemma 5.7. Then using Vietoris Begle mapping theorem from [19] we get a basis $\{p\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{m}$ for $H_{T^{n+1}}^{*}(Pl(k,n);\mathbb{Q})$ as a $\mathbb{Q}[y_{1},y_{2},\dots,y_{n},z]$ module, where $p\widetilde{\mathbb{S}}_{\lambda^{i}}$ is the isomorphic image of $\widetilde{\mathbb{S}}_{\lambda^{i}}$. Then we deduce that ${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}$ is the isomorphic image of $p\widetilde{\mathbb{S}}_{\lambda^{i}}$ after evaluating at $z=0$. We discuss the equivariant Pieri rule in $H_{T^{n}}^{*}(\mbox{Gr}_{\bf b}(k,n);\mathbb{Z})$. Moreover we compute the equivariant structure constants with respect to the basis $\{{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{m}$.
Theorem C (Theorem 5.14).
The equivariant structure constant ${\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{\ell}$ is given by
${\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{\ell}=\sum_{\lambda^{\ell}\succeq\lambda^{q}\succeq\lambda^{i},\lambda^{j}}\sum_{P:\partial P=\Delta_{\sigma_{r}\lambda^{i}~{}\sigma_{r}\lambda^{j}}^{\sigma_{r}\lambda^{q}}}\sum_{s=0}^{|P|}a^{(P)}_{s}{\bf{b}}\widetilde{\mathscr{C}}_{q~{}g^{s}}^{\ell}.$
Here $P$ is a puzzle and and $\delta P$ is the boundary of the the labeled equilateral triangle induced from $P$. The coefficients $a^{(P)}_{s}$ and ${\bf{b}}\widetilde{\mathscr{C}}_{q~{}g^{s}}^{\ell}$ can be computed using the equvariant structure cnstants $\widetilde{\mathscr{C}}_{i~{}j}^{\ell}$ in $H_{T^{n}}^{*}(\mbox{Gr}(k,n);\mathbb{Z})$ and the equivariant Pieri rule in $H_{T^{n}}^{*}(\mbox{Gr}_{\bf b}(k,n);\mathbb{Z})$ respectively, see (5.6) and (5.13). We demonstrate the positivity of equivariant structure constants; see Theorem 5.16. Using the upper triangularity of every basis element ${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}$, we prove that:
Theorem D (Theorem 5.15).
Let ${\rm Gr}_{\bf{b}}(k,n)$ be a divisive weighted Grassmann orbifold. Then all the equivariant structure constants ${\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{\ell}$ are in $\mathbb{Z}[y_{1},y_{2},\dots,y_{n}]$.
In Section 6, we work out on a basis $\{{\bf b}S_{\lambda^{i}}\}_{i=0}^{m}$ for the integral cohomology groups of divisive weighted Grassmann orbifolds. Finally we compute the structure constant ${\bf{b}}\mathscr{C}_{i~{}j}^{\ell}$ with respect to the above basis with integer coefficients:
Theorem E (Theorem 6.3).
The structure constant ${\bf{b}}\mathscr{C}_{i~{}j}^{\ell}$ is given by
${\bf{b}}\mathscr{C}_{i~{}j}^{\ell}=\mathscr{C}_{i~{}j}^{\ell}+\sum_{\lambda^{q}~{}:~{}\lambda^{\ell}\succ\lambda^{q}\succeq\lambda^{i},\lambda^{j}}\mathscr{D}_{i~{}j}^{\ell~{}q}.$
Here $\mathscr{C}_{i~{}j}^{\ell}$ is the structure constant in $H^{*}(\mbox{Gr}(k,n);\mathbb{Z})$ and $\mathscr{D}_{i~{}j}^{\ell~{}q}$ is described in (6.2). This describes the integral cohomology ring of a divisive $\mbox{Gr}_{\bf{b}}(k,n)$ explicitly. In particular, we describe the explicit computation of the structure constants and the integral cohomology ring of a divisive $\mbox{Gr}_{\bf{b}}(2,4)$.
2. Building sequences of weighted Grassmann orbifolds
In this Section, we introduce the notion of Plücker weight vector, which we use to give a new definition of weighted Grassmann orbifold. Then, we show that this definition includes the weighted Grassmannian discussed in [1] as well as [7]. We discuss a building sequence of $\mbox{Gr}_{\bf{b}}(k,n)$ that arises from a $q$-CW complex structure.
For two positive integers $k$ and $n$ such that $k<n$, we recall a Schubert symbol $\lambda$ is a sequence of $k$ integers $(\lambda_{1},\lambda_{2},\dots,\lambda_{k})$ such that $1\leq\lambda_{1}<\lambda_{2}<\cdots<\lambda_{k}\leq n$. A pair $(s,s^{\prime})$ is called a reversal of $\lambda$ if $s\in\lambda,s^{\prime}\notin\lambda$ and $s>s^{\prime}$. The set of all reversals of $\lambda$ is denoted by $\rm{rev}(\lambda)$. If $(s,s^{\prime})\in\rm{rev}(\lambda)$, then $(s,s^{\prime})\lambda$ denote the Schubert symbol obtained by replacing $s$ by $s^{\prime}$ in $\lambda$ and ordering the later set.
Define a partial order $\preceq$’ on the set of Schubert symbols for $k<n$ by
(2.1) $\lambda\preceq\mu\text{ if }\lambda_{i}\leq\mu_{i}\text{ for all }i=1,2,\dots,k.$
Then the Schubert symbols for $k<n$ form a lattice with respect to this partial order $\preceq$’.
There are ($n\choose k$ many Schubert symbols for two positive integers $k$ and $n$ such that $k<n$. Let $m={n\choose k}-1$. Now consider a total order
(2.2) $\lambda^{0}<\lambda^{1}<\lambda^{2}<\dots<\lambda^{m}$
on the set of all Schubert symbols which preserve the partial order $\preceq$’ i.e., if $\lambda^{i}\preceq\lambda^{j}$ then $i<j$. For example, one can consider the dictionary order on the set of Schubert symbols. Then, it is a total order that preserves the partial order $\preceq$ defined in (2.1).
2.1. Plücker weight vectors and weighted Grassmann orbifolds
Let $\Lambda^{k}(\mathbb{C}^{n})$ be the $k$-th exterior product of the complex $n$-dim space $\mathbb{C}^{n}$. Note that the standard ordered basis $\{e_{1},e_{2},\dots,e_{n}\}$ for $\mathbb{C}^{n}$ induces an ordered basis $\{e_{\lambda^{0}},e_{\lambda^{1}},\dots,e_{\lambda^{m}}\}$ of $\Lambda^{k}(\mathbb{C}^{n})$, where $e_{\lambda^{i}}=e_{\lambda^{i}_{1}}\wedge e_{\lambda^{i}_{2}}\wedge\dots\wedge e_{\lambda^{i}_{k}}$ for a schubert symbol $\lambda^{i}=(\lambda^{i}_{1},\lambda^{i}_{2},\dots,\lambda^{i}_{k})$. Then one can identify $\Lambda^{k}(\mathbb{C}^{n})$ with $\mathbb{C}^{m+1}$. Consider a subset of $\Lambda^{k}(\mathbb{C}^{n})$ defined by the following
(2.3) $Pl(k,n):=\{a_{1}\wedge a_{2}\wedge a_{3}\wedge\dots\wedge a_{k}\colon a_{i}\in\mathbb{C}^{n}\}\setminus\{\bf{0}\}.$
Then $Pl(k,n)\subseteq\mathbb{C}^{m+1}\setminus\{\bf{0}\}$. Consider an element $z=\sum_{i=0}^{m}z_{i}e_{\lambda^{i}}\in\mathbb{C}^{m+1}\setminus\{\bf{0}\}$. Then $z\in Pl(k,n)$ iff it satisfies the following condition
(2.4) $\sum_{j=0}^{k}(-1)^{j}z_{r_{j}}z_{s_{j}}=0,$
where $\lambda^{r_{j}}=(i_{1},...,i_{k-1},\ell_{j})$ and $\lambda^{s_{j}}=(\ell_{0},...,\widehat{\ell}_{j},...,\ell_{k})$ for any two ordered sequences $1\leq i_{1}<\dots<i_{k-1}\leq n$ and $1\leq\ell_{0}<\ell_{1}<\dots<\ell_{k}\leq n$, see [11, Theorem 3.4.11].
The relation described in (2.4) is known as a Plücker relation. The coordinates of $Pl(k,n)$ are known as Plücker coordinates.
Remark 2.1.
In (2.4), we can assume $2k\leq n$. If $2k>n$ consider $k^{\prime}=n-k$, then $2k^{\prime}=2n-2k<n$. Note that (${n\choose k}={n\choose k^{\prime}}$. Let $\lambda=(\lambda_{1},\dots,\lambda_{k})$ be a Schubert symbol for $k<n$. Then $\{1,2,\dots,n\}\setminus\{\lambda_{1},\dots,\lambda_{k}\}$ in a increasing order gives a Schubert symbol for $k^{\prime}<n$. This correspondence of the Schubert symbol gives a relation as in (2.4) in $Pl(k,n)$ induced from $Pl(k^{\prime},n)$.
Consider the ${\bf{b}}$-action of $\mathbb{C}^{*}$ on $\mathbb{C}^{m+1}\setminus\{{\bf 0}\}$ defined in (1.1). Now the natural question is whether $Pl(k,n)$ is a $\mathbb{C}^{*}$-invariant subset of $\mathbb{C}^{m+1}\setminus\{{\bf 0}\}$ under the ${\bf{b}}$-action defined in (1.1). This is not true in general for any arbitrary ${\bf{b}}=(b_{0},b_{1},\dots,b_{m})\in(\mathbb{Z}_{\geq 1})^{m+1}$. Note that $Pl(k,n)$ will be an invariant subset of $\mathbb{C}^{m+1}\setminus\{{\bf 0}\}$ iff
$(t^{b_{0}}z_{0},t^{b_{1}}z_{1},,\dots,t^{b_{m}}z_{m})\in Pl(k,n)\text{ for every }t\in\mathbb{C}^{*}\text{ and }(z_{0},z_{1},\dots,z_{m})\in Pl(k,n).$
Using (2.4), $(t^{b_{0}}z_{0},t^{b_{1}}z_{1},\dots,t^{b_{m}}z_{m})\in Pl(k,n)$ iff
(2.5) $\sum_{j=0}^{k}(-1)^{j}t^{b_{r_{j}}+b_{s_{j}}}z_{r_{j}}z_{s_{j}}=0.$
Remark 2.2.
Consider two ordered sequences $1\leq i_{1}<\dots<i_{k-1}\leq n$ and $1\leq\ell_{0}<\ell_{1}<\dots<\ell_{k}\leq n$. Then we have $k+1$ pairs $(r_{0},s_{0}),\dots,(r_{k},s_{k})$. Consider two pair $(r_{i},s_{i})$, $(r_{j},s_{j})$ for $i\neq j$ and define $z=(z_{0},z_{1},\dots,z_{m})$ by $z_{r_{i}}=z_{s_{i}}=z_{r_{j}}=1$ and $z_{s_{j}}=(-1)^{i+j+1}$. Then $z\in Pl(k,n)$ and (2.5) reduced to $t^{b_{r_{j}}+b_{s_{j}}}=t^{b_{r_{i}}+b_{s_{i}}}$, for all $t\in\mathbb{C}^{*}$. This implies $b_{r_{i}}+b_{s_{i}}=b_{r_{j}}+b_{s_{j}}$. Thus both (2.4) and (2.5) satisfies together for any $t\in\mathbb{C}^{*}$ and for every $z\in Pl(k,n)$ iff $b_{r_{j}}+b_{s_{j}}$ is constant for every $j=0,1,\dots,k$. This motivates us to introduce the Plücker weight vector.
Note that a Plücker relation $\mathfrak{p}$ as in (2.4) uniquely corresponds to $(k+1)$ many pairs $\{(r_{0}^{\mathfrak{p}},s_{0}^{\mathfrak{p}}),(r_{1}^{\mathfrak{p}},s_{1}^{\mathfrak{p}}),\dots,(r_{k}^{\mathfrak{p}},s_{k}^{\mathfrak{p}})\}$. In the next definition and throughout the whole paper, we assume $k$ and $n$ are two positive integers with $k<n$ and $m+1={n\choose k}$.
Definition 2.3.
A weight vector ${\bf{b}}=(b_{0},b_{1},\dots,b_{m})\in(\mathbb{Z}_{\geq 1})^{m+1}$ is called Plücker weight vector if for every Plücker relation $\mathfrak{p}$
[…]
for all $j\in\{0,1,\dots,k\}$, where the $(k+1)$ many pairs $\{(r_{0}^{\mathfrak{p}},s_{0}^{\mathfrak{p}}),(r_{1}^{\mathfrak{p}},s_{1}^{\mathfrak{p}}),\dots,(r_{k}^{\mathfrak{p}},s_{k}^{\mathfrak{p}})\}$ are uniquely defined for every Plücker relation $\mathfrak{p}$.
Remark 2.4.
Let ${\bf b}=(b_{0},b_{1},\dots,b_{m})$ be a Plücker weight vector. If $\lambda^{i}\cup\lambda^{j}=\lambda^{r}\cup\lambda^{q}$ for any 4 Schubert symbols $\lambda^{i}$, $\lambda^{j}$, $\lambda^{q}$, $\lambda^{r}$ then $b_{i}+b_{j}=b_{r}+b_{q}$.
Example 2.5.
For $k=2$ and $n=4$, we have $m={4\choose 2}-1=5$. In this case, we have $6$ Schubert symbols which are
$\lambda^{0}=(1,2),\lambda^{1}=(1,3),\lambda^{2}=(1,4),\lambda^{3}=(2,3),\lambda^{4}=(2,4)~{}\text{and}~{}\lambda^{5}=(3,4)$
in the ordering as in (2.2). Consider $z=(z_{0},z_{1},\dots,z_{5})\in\mathbb{C}^{6}\setminus\{0\}$. Let
$i_{1}=1,~{}\ell_{0}=2<\ell_{1}=3<\ell_{2}=4.$
Then $z\in Pl(2,4)$ iff $z_{0}z_{5}-z_{1}z_{4}+z_{2}z_{3}=0$. A weight vector ${\bf{b}}=(b_{0},b_{1},\dots,b_{5})\in(\mathbb{Z}_{\geq 1})^{6}$ is a Plücker weight vector if $b_{0}+b_{5}=b_{1}+b_{4}=b_{2}+b_{3}$.
Using Remark 2.2, it follows that $Pl(k,n)$ is $\mathbb{C}^{*}$-invariant subset of $\mathbb{C}^{m+1}\setminus\{{\bf 0}\}$ under the ${\bf b}$-action iff ${\bf b}$ is a Plücker weight vector. This paves the way to define the following.
Definition 2.6.
Let $k$ and $n$ be two positive integers such that $k<n$ and $m+1={n\choose k}$. If ${\bf{b}}=(b_{0},b_{1},\dots,b_{m})\in(\mathbb{Z}_{\geq 1})^{m+1}$ is a Plücker weight vector there is a ${\bf b}$-action of $\mathbb{C}^{*}$ on $Pl(k,n)$ as in (1.1). We denote the orbit space
$\mbox{Gr}_{\bf{b}}(k,n):=\frac{Pl(k,n)}{{\bf b}\text{-action}}.$
Consider the quotient map
$\pi_{\bf{b}}:Pl(k,n)\to\mbox{Gr}_{\bf{b}}(k,n)$
defined by $\pi_{\bf{b}}(z)=[z]$, where $[z]$ is the orbit of $z$ with respect to the ${\bf b}$-action. The topology on $\mbox{Gr}_{\bf{b}}(k,n)$ is given by the quotient topology induced by the map $\pi_{\bf{b}}$.
Remark 2.7.
If ${\bf{b}}=(1,1,\dots,1)$ then ${\bf{b}}$ is trivially a Plücker weight vector. In this case the space $\mbox{Gr}_{\bf{b}}(k,n)$ is the Grassmann manifold $\mbox{Gr}(k,n)$. We denote the corresponding quotient map by
$\pi:Pl(k,n)\to\mbox{Gr}(k,n).$
Now we recall the definition of weighted Grassmannians from [1] and show that the space $\mbox{Gr}_{\bf b}(k,n)$ in Definition 2.6 is a generalization of $\mbox{WGr}(k,n)$ in [1, Definition 2.2].
Let $W:=(w_{1},w_{2},\dots,w_{n})\in(\mathbb{Z}_{\geq 0})^{n}$ and $a\in\mathbb{Z}_{\geq 1}$. For $k<n$, recall $Pl(k,n)$ from (2.3). The algebraic torus (*$(\mathbb{C}^{*})^{n+1}$ acts on $Pl(k,n)$ by the following
(2.6) $(t_{1},t_{2},\dots,t_{n},t)\sum_{i=0}^{m}z_{i}e_{\lambda^{i}}=\sum_{i=0}^{m}t{\cdot}t_{\lambda^{i}}z_{i}e_{\lambda^{i}}$
Consider a subgroup $WD$ of (*$(\mathbb{C}^{*})^{n+1}$ defined by
(2.7) $WD:=\{(t^{w_{1}},t^{w_{2}},\dots,t^{w_{n}},t^{a}):t\in\mathbb{C}^{*}\}.$
Then there is a restricted action of $WD$ on $Pl(k,n)$. Abe and Matsumura called the quotient $\frac{Pl(k,n)}{WD}$ a weighted Grassmannian and denote it by $\mbox{WGr}(k,n)$ in [1].
For all $i\in\{0,1,\dots,m\}$, let
(2.8) $b_{i}=w_{\lambda^{i}}:=a+\sum_{j=1}^{k}w_{\lambda_{j}^{i}},$
where $\lambda^{i}=(\lambda_{1}^{i},\lambda_{2}^{i},\dots,\lambda_{k}^{i})$ is the $i$-th Schubert symbol given in (2.2). Then $b_{i}\geq 1$ for all $i\in\{0,1,\dots,m\}$. Consider the weight vector ${\bf{b}}=(b_{0},b_{1},\dots,b_{m})\in(\mathbb{Z}_{\geq 1})^{m+1}$. From the Definition 2.3 it follows that ${\bf{b}}$ is a Plücker weight vector. Note that action of $WD$ on $Pl(k,n)$ is same as the ${\bf{b}}$-action of $\mathbb{C}^{*}$ on $Pl(k,n)$. Thus $\mbox{Gr}_{\bf{b}}(k,n)$ is same as the weighted Grassmannian $\mbox{WGr}(k,n)$, where the Plücker weight vector ${\bf{b}}=(b_{0},\dots,b_{m})$ is defined using (2.8).
Now, we check whether the converse is true. More explicitly, for a Plücker weight vector ${\bf{b}}=(b_{0},\dots,b_{m})\in(\mathbb{Z}_{\geq 1})^{m+1}$ whether there exist $W:=(w_{1},w_{2},\dots,w_{n})\in(\mathbb{Z}_{\geq 0})^{n}$ and $a\in\mathbb{Z}_{\geq 1}$ such that $b_{i}=w_{\lambda^{i}}$.
Remark 2.8.
For a given Plücker weight vector ${\bf{b}}=(b_{0},b_{1},\dots,b_{m})\in(\mathbb{Z}_{\geq 1})^{m+1}$, where $m+1={n\choose k}$ for $k<n$, there may not exist $W=(w_{1},w_{2}\dots,w_{n})\in(\mathbb{Z}_{\geq 0})^{n}$ and $a\in\mathbb{Z}_{\geq 1}$ such that $b_{i}=w_{\lambda^{i}}$. Let $n=4$ and $k=2$. Then $m={4\choose 2}-1=5$. In this case, we have $6$ Schubert symbols given by
$\lambda^{0}=(1,2)<\lambda^{1}=(1,3)<\lambda^{2}=(1,4)<\lambda^{3}=(2,3)<\lambda^{4}=(2,4)<\lambda^{5}=(3,4).$
Consider a weight vector ${\bf{b}}=(5,1,4,3,6,2)\in(\mathbb{Z}_{\geq 1})^{6}$. Then ${\bf{b}}$ is a Plücker weight vector follows from Example 2.5. If there exist $W=(w_{1},w_{2},w_{3},w_{4})\in(\mathbb{Z}_{\geq 0})^{4}$ and $a\in\mathbb{Z}_{\geq 1}$ such that $b_{i}=w_{\lambda^{i}}$ then $b_{1}=w_{1}+w_{3}+a=1$. Now $W\in(\mathbb{Z}_{\geq 0})^{4}$ implies that $b_{1}\geq a$. The only choice for $a$ is $1$. Then we have $w_{1}+w_{3}=0$. Also $w_{1}-w_{3}=b_{0}-b_{3}=2$. Which gives $w_{1}=1$, $w_{3}=-1$, $w_{2}=3$ and $w_{4}=2$. Note that here $W\notin(\mathbb{Z}_{\geq 0})^{4}$.
If we choose another Plücker weight vector ${\bf{b}}=(5,1,4,4,7,3)$, then only choice of $a=1$. Similar to the above calculation we get $W=(1/2,7/2,-1/2,5/2)$.
In the following Proposition, we prove that for a Plücker weight vector ${\bf b}=(b_{0},\dots,b_{m})$, there exist solutions of $W$ and $a$ in a broader class.
Proposition 2.9.
Let ${\bf{b}}=(b_{0},b_{1},\dots,b_{m})\in(\mathbb{Z}_{\geq 1})^{m+1}$ be a Plücker weight vector for $k<n$. Then there exist $a\in\{1,2,\dots,k\}$ and $W=(w_{1},w_{2},\dots,w_{n})\in\mathbb{Z}^{n}$ such that $b_{i}=w_{\lambda^{i}}$ for all $i=0,1,\dots,m$, where $w_{\lambda^{i}}$ is defined in (2.8).
Proof.
Consider the linear equations defined by $w_{\lambda^{i}}=b_{i}$ for all $i=0,1,\dots,m$ with variables $w_{1},w_{2},\dots,w_{n}\text{ and }a$. Let us fix the variable $w_{1}$. For every $j$ satisfying $2\leq j\leq n$ we can choose two Schubert symbols $\lambda^{r_{j}}$ and $\lambda^{s_{j}}$ such that $(1,j)\lambda^{r_{j}}=\lambda^{s_{j}}$. Then we have linear equations of the form
(2.9) $w_{1}-w_{j}=b_{r_{j}}-b_{s_{j}};\text{ for }2\leq j\leq n.$
Note that if $(1,j)\lambda^{q_{j}}=\lambda^{t_{j}}$ for any two Schubert symbols $\lambda^{q_{j}}$ and $\lambda^{t_{j}}$ then $\lambda^{r_{j}}\cap\lambda^{t_{j}}=\lambda^{q_{j}}\cap\lambda^{s_{j}}$. Since ${\bf b}$ is a Plücker weight vector, using Remark 2.4, $b_{r_{j}}+b_{t_{j}}=b_{q_{j}}+b_{s_{j}}$. So, $b_{r_{j}}-b_{s_{j}}=b_{q_{j}}-b_{t_{j}}$. Thus (2.9) is well defined if ${\bf b}$ is a Plücker weight vector.
Now fix a Schubert symbol $\lambda^{i}=(2,3,\dots,k+1)$. We add the $k$ equations of (2.9) for $12\leq j\leq k+$ 1 with the equation $w_{\lambda^{i}}=b_{i}$ and we get
$kw_{1}+a=b_{i}+\sum_{j=2}^{k+1}(b_{r_{j}}-b_{s_{j}}).$
Note that the right side of the above equation is an integer. This gives a solution for $a\in\{1,2,\dots,k\}$ which satisfies
$b_{i}+\sum_{j=2}^{k+1}(b_{r_{j}}-b_{s_{j}})\equiv a~{}(\text{mod }k).$
Thus $kw_{1}\equiv 0~{}(\text{mod }k)$. Then, we get an integer solution for $w_{1}$. Consequently, we get integer solutions for $w_{j}$ for all $2\leq j\leq n$ using (2.9). Thus all the $w_{j}$ are integers. ∎
Remark 2.10.
For a given Plücker weight vector ${\bf{b}}=(b_{0},\dots,b_{m})\in(\mathbb{Z}_{\geq 1})^{n}$ there are infinitely many solutions of the form $W^{\prime}=(w_{1}-\alpha,\dots,w_{n}-\alpha)\in\mathbb{Z}^{n}$ and $a^{\prime}=a+k\alpha$ for any $\alpha\in\mathbb{Z}$ such that $b_{i}=w_{\lambda^{i}}$.
Thus, Definition 2.6 broadens the class of weighted Grassmannian in [1] as well as [7] in terms of Plücker weight vector. The authors in [1] discussed the orbifold structure of $\mbox{WGr}(k,n)$ for $W=(w_{1},\dots,w_{n})\in(\mathbb{Z}_{\geq 0})^{n}$ and $a\in\mathbb{Z}_{\geq 1}$. Using the argument of [1, Subsection 2.2] and Proposition 2.9, the quotient space $\mbox{Gr}_{\bf{b}}(k,n)$ has an orbifold structure for a Plücker weight vector ${\bf{b}}$. We call the space $\mbox{Gr}_{\bf{b}}(k,n)$ a weighted Grassmann orbifold associated to the Plücker weight vector ${\bf{b}}$.
2.2. $q$-CW complex structures of weighted Grassmann orbifolds
In this subsection, we discuss the $q$-CW complex structure of $\mbox{Gr}_{\bf{b}}(k,n)$ following [7, 1]. This $q$-CW complex structure induces a building sequence on $\mbox{Gr}_{\bf b}(k,n)$. Then, we explore the cofibrations that appear at each stage of the building sequence.
The CW complex structure of $\mbox{Gr}(k,n)$ given by
$\mbox{Gr}(k,n)=\bigsqcup_{i=0}^{m}{E}(\lambda^{i}),\mbox{ where }$
${E}(\lambda^{i})=\{[(z_{0},z_{1},\dots,z_{m})]\in\mbox{Gr}(k,n)~{}|~{}z_{i}=1,z_{j}=0{\mbox{ for }}j>i\}.$
Every element of $E(\lambda^{i})$ has a presentation $[A]=[a_{1}\wedge a_{2}\wedge\dots\wedge a_{k}]$. If $\lambda^{i}=(\lambda^{i}_{1},\lambda^{i}_{2},\dots,\lambda^{i}_{k})$ is the $i$-th Schubert symbol then
(2.10) as:=($a_{s}:=(a_{s1},a_{s2},\dots,a_{s{\lambda^{i}_{s}}},\dots,0)\in\mathbb{C}^{n},$
such that ${a}_{s\lambda^{i}_{s}}=1$ and ${a}_{s\lambda^{i}_{j}}=0$ for $11\leq j\leq s-$ 1 and $1\leq s\leq k$, see [16].
Then $E(\lambda^{i})$ is an open cell of $\mbox{Gr}(k,n)$ of dimension $2d(\lambda^{i})$, where
$d(\lambda^{i}):=(\lambda^{i}_{1}-1)+(\lambda^{i}_{2}-2)+\cdots+(\lambda^{i}_{k}-k).$
For every $i\in\{0,1,\dots,m\}$ consider a subset of Schubert symbol
(2.11) $R(\lambda^{i})=\{\lambda^{j}:\lambda^{j}=(\lambda^{i}_{s},r)\lambda^{i}\text{ for }(\lambda^{i}_{s},r)\in{\rm{rev}}(\lambda^{i})\text{ and }s=1,2,\dots,k\}.$
Note that $|R(\lambda^{i})|=d(\lambda^{i})\leq i$ for all $i=0,1,\dots,m$. Consider $\mathbb{C}^{d(\lambda^{i})}\subset\mathbb{C}^{m+1}$ by
$\mathbb{C}^{d(\lambda^{i})}=\{(z_{0},z_{1},\dots,z_{m})~{}|~{}z_{j}=0\text{ if}\lambda^{j}\notin R(\lambda^{i})\}.$
Then, ${E}(\lambda^{i})$ is homeomorphic to $\mathbb{C}^{d(\lambda^{i})}$ by the following map
$g\colon{E}(\lambda^{i})\to\mathbb{C}^{d(\lambda^{i})}.$
The $j$-th entry of $g([a_{1}\wedge\dots\wedge a_{k}])$ is equal to $a_{sr}$ (the $r$-th entry of $a_{s}$) if $\lambda^{j}=(\lambda^{i}_{s},r)\lambda^{i}$ for some $1\leq s\leq k$, $1\leq r\leq n$ and $r\notin\{\lambda^{i}_{1},\dots,\lambda^{i}_{k}\}$. This defines both $g$ and $g^{-1}$ completely. Note that this is a bijective correspondence, and both $g$ and $g^{-1}$ are continuous.
Let $\widetilde{E}(\lambda^{i})=\pi^{-1}(E(\lambda^{i}))$. Then
$\widetilde{E}(\lambda^{i})=\{(z_{0},z_{1},\dots,z_{m})\in Pl(k,n)~{}|~{}z_{i}\neq 0,z_{j}=0\text{ for }j>i\}.$
Note that $\widetilde{E}(\lambda^{i})$ is homeomorphic to ${E}(\lambda^{i})\times\mathbb{C}^{*}$ via the map
$f:\widetilde{E}(\lambda^{i})\to{E}(\lambda^{i})\times\mathbb{C}^{*}\text{ defined by }f(z)=([z],z_{i}).$
The inverse map is defined by
${E}(\lambda^{i})\times\mathbb{C}^{*}\to\widetilde{E}(\lambda^{i});~{}~{}([z],s)\to s(z_{i})^{-1}z.$
Now $\widetilde{E}(\lambda^{i})\subset Pl(k,n)$ and there is a there is a restricted $\bf b$-action of $\mathbb{C}^{*}$ on $\widetilde{E}(\lambda^{i})$. Define a ${\bf{b}}$-action on ${E}(\lambda^{i})\times\mathbb{C}^{*}$ by
$t([z],s)=([tz],t^{b_{i}}s).$
Then $f$ is $\mathbb{C}^{*}$-equivariant with respect to the above ${\bf{b}}$-actions.
Note that $g$ is also $\mathbb{C}^{*}$-equivariant map, where the action of $\mathbb{C}^{*}$ on $\mathbb{C}^{d(\lambda^{i})}$ is restriction of the ${\bf{b}}$-action on $\mathbb{C}^{m+1}$. Then
$(g,\mbox{id})\circ f:\widetilde{E}(\lambda^{i})\to\mathbb{C}^{d(\lambda^{i})}\times\mathbb{C}^{*}$
is an $\mathbb{C}^{*}$-equivariant homeomorphism. Let $G(b_{i})$ be the group of $b_{i}$-th roots of unity. The finite subgroup $G(b_{i})$ of $\mathbb{C}^{*}$ acts on the second factor trivially. Thus
$\pi_{\bf{b}}(\widetilde{E}(\lambda^{i}))\cong\frac{\mathbb{C}^{d(\lambda^{i})}\times\mathbb{C}^{*}}{{\bf b}\text{-action}}\cong\frac{\mathbb{C}^{d(\lambda^{i})}}{G(b_{i})}.$
Thus the $q$-cell decomposition of $\mbox{Gr}_{\bf b}(k,n)$ is given by $\mbox{Gr}_{\bf{b}}(k,n)=\sqcup_{i=0}^{m}\frac{\mathbb{C}^{d(\lambda^{i})}}{G(b_{i})}.$
For each $j\in\{0,1,\dots,m\}$, define $X_{j}=\sqcup_{i=0}^{j}\frac{\mathbb{C}^{d(\lambda^{i})}}{G(b_{i})}.$ The building sequence corresponding to the $q$-CW complex structure is
(2.12) $\{pt\}=X^{0}\subset X^{1}\subset X^{2}\subset\dots\subset X^{m}=\mbox{Gr}_{\bf b}(k,n).$
Next, we recall lens complex from [13]. Let $(c_{0},c_{1},\dots,c_{m})\in(\mathbb{Z}_{\geq 1})^{m+1}$ and $G(c_{m})$ be the $c_{m}$-th roots of unity.
Definition 2.11.
[13] Consider an action of $G(c_{m})$ on $S^{2m-1}$ defined by
$t(z_{0},z_{1},\dots,z_{m-1})=(t^{c_{0}}z_{0},t^{c_{1}}z_{1},\dots,t^{c_{m-1}}z_{m-1}),$
where $t\in G(c_{m})$. The quotient space
$L^{\prime}(c_{m};(c_{0},c_{1},\dots,c_{m-1})):=\dfrac{S^{2m-1}}{G(c_{m})}$
is called the lens complex. If $c_{0},\dots,c_{m-1}$ are prime to $c_{m}$ then it is a usual generalized lens space.
Now consider a Plücker weight vector ${\bf b}=(b_{0},b_{1},\dots,b_{m})$. For every $i=1,2,\dots,m$ define a weight vector
(2.13) ${\bf{b}}^{(i)}:=\{(b_{0},b_{1},\dots,b_{i-1})\in(\mathbb{Z}_{\geq 1})^{d(\lambda^{i})}\text{ such that }b_{j}=0\text{ if }\lambda^{j}\notin R(\lambda^{i})\}.$
We denote the following lens complex by $L^{\prime}(b_{i};{\bf{b}}^{(i)})$
$L^{\prime}(b_{i};{\bf{b}}^{(i)}):=\{[z_{0},z_{1},\dots,z_{i-1}]\in L^{\prime}(b_{i};(b_{0},b_{1},\ldots,b_{i-1}))~{}|~{}z_{j}=0\text{ if }\lambda^{j}\notin R(\lambda^{i})\}$
Proposition 2.12.
$L^{\prime}(b_{i};{\bf{b}}^{(i)})\hookrightarrow X^{i-1}\to X^{i}$ is a cofibration for every $i=1,\dots,m$.
Proof.
Consider $\mathbb{C}^{d(\lambda^{i})}=\{(z_{0},z_{1},\dots,z_{m})~{}|~{}z_{j}=0\text{ if}\lambda^{j}\notin R(\lambda^{i})\}\subset\mathbb{C}^{m}$ and denote the unit circle on $\mathbb{C}^{d(\lambda^{i})}$ by $S^{2d(\lambda^{i})-1}=\{z\in\mathbb{C}^{d(\lambda^{i})}~{}|~{}|z|=1\}$. Then
$L^{\prime}(b_{i};{\bf{b}}^{(i)})=\dfrac{S^{2d(\lambda^{i})-1}}{G(b_{i})}.$
Using the $q$-cell decomposition $X^{i}=X^{i-1}\sqcup\frac{\mathbb{C}^{d(\lambda^{i})}}{G(b_{i})}$. Now the result follows from the fact that $\frac{\mathbb{C}^{d(\lambda^{i})}}{G(b_{i})}$ is an unbounded cone over $L^{\prime}(b_{i};{\bf{b}}^{(i)})$. ∎
3. Classification of weighted Grassmann orbifolds upto certain homeomorphism
In this section, we introduce the Plücker permutation that induces a permutation on Plücker coordinates up to some sign. Then we prove that a Plücker permutation induces an equivariant homeomorphism between two weighted Grassmann orbifolds. We also show that two weighted Grassmann orbifolds are homeomorphic if their Plücker weight vectors are differed by scalar multiplication. Conversely, we prove that if two weighted Grassmann orbifolds are homeomorphic with a small assumption, their weight vectors differ by scalar multiplication and a permutation. This provides an equivariant classification of a weighted Grassmann orbifold in terms of Plücker weight vectors.
Let $\sigma$ be a permutation on $\{0,1,2,\dots,m\}$ and $z=(z_{0},\dots,z_{m})\in\mathbb{C}^{m+1}\setminus\{{\bf 0}\}$. Define $\sigma z:=(z_{\sigma(0)},z_{\sigma(1)},\dots,z_{\sigma(m)})$. Consider a sign vector $t=(t_{0},\dots,t_{m})\in\{1,-1\}^{m+1}$, i.e., $t_{i}\in\{1,-1\}$ for all $i=0,1,\dots,m$. Define $t{\cdot}\sigma z:=(t_{0}z_{\sigma(0)},t_{1}z_{\sigma(1)},\dots,t_{m}z_{\sigma(m)})$.
Definition 3.1.
A permutation $\sigma$ on $\{0,1,2,\dots,m\}$ is said to be Plücker permutation if there exist $t_{\sigma}\in\{1,-1\}^{m+1}$ such that $t_{\sigma}{\cdot}\sigma z\in Pl(k,n)$ for every $z\in Pl(k,n)$.
Let $\sigma$ and $\eta$ be two Plücker permutations on $\{0,1,2,\dots,m\}$. Then $\sigma\circ\eta$ is also a Plücker permutation on $\{0,1,2,\dots,m\}$. Note that, $\sigma^{-1}=\sigma^{c}$ for some positive integer $c$. Thus $\sigma^{-1}$ is also a Plücker permutation on $\{0,1,2,\dots,m\}$. Hence the Plücker permutations on $\{0,1,\dots,m\}$ form a subgroup of the permutation group on $\{0,1,\dots,m\}$.
Remark 3.2.
  1. (1)
    Let $\lambda^{i}=(\lambda_{1}^{i},\ldots,\lambda_{k}^{i})$ be the $i$-th Schubert symbol described in (2.2), for $i=0,\ldots,m$ and $\phi\in S_{n}$. Define $\phi\lambda^{i}:=(\phi(\lambda_{i_{1}}^{i}),\ldots,\phi(\lambda_{i_{k}}^{i}))$, where $i_{1},\dots,i_{k}\in\{1,\dots,k\}$ such that $1\leq\phi(\lambda_{i_{1}}^{i})<\phi(\lambda_{i_{2}}^{i})<\cdots<\phi(\lambda_{i_{k}}^{i})\leq n$. Then $\phi\lambda^{i}\in\{\lambda^{0},\dots,\lambda^{m}\}$. Thus $\phi\in S_{n}$ induces a permutation $\sigma$ on $\{0,1,\dots,m\}$ defined by $\phi\lambda^{i}=\lambda^{\sigma(i)}$. Then $\sigma$ is a Plücker permutation on $\{0,1,\dots,m\}$ follows from (2.4). Thus there exists an injective group homomorphism from $S_{n}$ to the group of the Plücker permutations on $\{0,1,\dots,m\}$ if $m+1={n\choose k}$.
  2. (2)
    The converse is not always true. For a particular case $n=4,k=2$, the total number of Plücker permutations are $48$. Among them, a total of $24$ permutations forms $S_{4}$.
Lemma 3.3.
Let $\sigma$ be a Plücker permutation. Then $\sigma{\bf{b}}$ is a Plücker weight vector for every Plücker weight vector ${\bf{b}}$.
Proof.
Let ${\bf{b}}=(b_{0},\dots,b_{m})$ be a Plücker weight vector. For every $(z_{0},\dots,z_{m})\in Pl(k,n)$ we have $(t^{b_{0}}z_{0},\dots,t^{b_{m}}z_{m})\in Pl(k,n)$ fot every $t\in\mathbb{C}^{*}$. If $\sigma$ is a Plücker permutation on $\{0,1,\dots,m\}$ then there exist $t_{\sigma}=(t_{0},\dots,t_{m})\in\{1,-1\}^{m+1}$ such that $(t_{0}z_{\sigma(0)},\dots,t_{m}z_{\sigma(m)})$ is in $Pl(k,n)$. Moreover, $t_{\sigma}{\cdot}\sigma(t^{b_{0}}z_{0},\dots,t^{b_{m}}z_{m})$ is also a Plücker coordinate. Thus $(t_{0}t^{b_{\sigma(0)}}z_{\sigma(0)},\dots,t_{m}t^{b_{\sigma(m)}}z_{\sigma(m)})\in Pl(k,n)$. Then using Remark 2.2, it follows that $\sigma{\bf{b}}=(b_{\sigma(0)},\dots,b_{\sigma(m)})$ is a Plücker weight vector. ∎
Now we can consider the weighted Grassmann orbifolds $\mbox{Gr}_{\sigma{\bf{b}}}(k,n)$ for a Plücker permutations $\sigma$. Plücker permutations play an important role in characterizing the weighted Grassmann orbifold up to a weakly equivariant homeomorphism. Consider the (*$(\mathbb{C}^{*})^{n}$-action on $\mbox{Gr}_{\bf{b}}(k,n)$ defined by
(3.1) $(t_{1},t_{2},\dots,t_{n})\sum_{i=0}^{m}[a_{i}e_{\lambda^{i}}]=\sum_{i=0}^{m}[t_{\lambda^{i}}a_{i}e_{\lambda^{i}}]$
where $t_{\lambda}:=t_{\lambda_{1}}\cdots t_{\lambda_{k}}$, for the Schubert symbol $\lambda=(\lambda_{1},\ldots,\lambda_{k})$, for $k<n$.
Theorem 3.4.
[…]
Proof.
A Plücker permutation $\sigma$ induces the natural weakly equivariant homeomorphism $\bar{f}_{\sigma}\colon Pl(k,n)\to Pl(k,n)$ defined by $\bar{f}_{\sigma}(z)=t_{\sigma}{\cdot}\sigma z$. Then $\bar{f}_{\sigma}$ induces a homeomorphism $f_{\sigma}$ between $\mbox{Gr}_{\bf{b}}(k,n)$ and $\mbox{Gr}_{\sigma{\bf{b}}}(k,n)$ such that the following diagram commutes.
[…]
The homeomorphism $f_{\sigma}$ becomes weakly equivariant with respect to the permuted (*$(\mathbb{C}^{*})^{n}$-action on $\mbox{Gr}_{\sigma{\bf b}}(k,n)$ defined by
(3.2) $(t_{1},t_{2},\dots,t_{n})\sum_{i=0}^{m}[a_{i}e_{\lambda^{i}}]=\sum_{i=0}^{m}[t_{\lambda^{\sigma(i)}}a_{i}e_{\lambda^{i}}]$
Now we discuss how the Plücker permutation $\sigma$ affects the $q$-CW complex structure on $\mbox{Gr}_{\bf{b}}(k,n)$. The permutation of the coordinates in $\mathbb{C}^{m+1}$ induces another decomposition of $Pl(k,n)$.
(3.3) $Pl(k,n)=\sqcup_{i=0}^{m}\sigma(\widetilde{E}(\lambda^{i})).$
Now $\sigma(\widetilde{E}(\lambda^{i}))$ is equivariantly homeomorphic to $\sigma\mathbb{C}^{d(\lambda^{i})}\times\mathbb{C}^{*}$, where the action of $\mathbb{C}^{*}$ on $\sigma(\widetilde{E}(\lambda^{i}))$ and $\sigma\mathbb{C}^{d(\lambda^{i})}$ are restriction of the ${\bf{b}}$-action on $\mathbb{C}^{m+1}$. The action of $\mathbb{C}^{*}$ on the $\sigma\mathbb{C}^{d(\lambda^{i})}\times\mathbb{C}^{*}$ is given by
$t(z,s)=(tz,t^{b_{\sigma(i)}}s).$
Thus we have the following
$\pi_{b}(\sigma(\widetilde{E}(\lambda^{i})))\cong\frac{\sigma(\widetilde{E}(\lambda^{i}))}{{\bf{b}}\mbox{-action }}\cong\frac{\sigma(\mathbb{C}^{d(\lambda^{i})})}{G(b_{\sigma(i)})}.$
So we get a $q$-CW complex structure of the weighted Grassmann orbifold $\mbox{Gr}_{\bf{b}}(k,n)$ given by
$\mbox{Gr}_{\bf{b}}(k,n)=\frac{\sigma(\mathbb{C}^{d(\lambda^{0})})}{G(b_{\sigma(0)})}\sqcup\frac{\sigma(\mathbb{C}^{d(\lambda^{1})})}{G(b_{\sigma(1)})}\sqcup\dots\sqcup\frac{\sigma(\mathbb{C}^{d(\lambda^{m})})}{G(b_{\sigma(m)})}.$
Also, the usual $q$-CW complex structure on a weighted Grassmann orbifold, as in Subsection 2.2, gives the following $q$-CW complex structure.
$\mbox{Gr}_{\sigma{\bf b}}(k,n)=\frac{\mathbb{C}^{d(\lambda^{0})}}{G(b_{\sigma(0)})}\sqcup\frac{\mathbb{C}^{d(\lambda^{1})}}{G(b_{\sigma(1)})}\sqcup\dots\sqcup\frac{\mathbb{C}^{d(\lambda^{m})}}{G(b_{\sigma(m)})}.$
Remark 3.5.
Let ${\bf{b}}=(b_{0},\dots,b_{m})$ be a Plücker weight vector corresponding to some $(W,a)$ and $\phi\in S_{n}$. Let $W^{\prime}:=(w_{\phi(1)},\dots,w_{\phi(n)})$ and $\{\phi b_{0},\dots,\phi b_{m}\}$ be the Plücker weight corresponding to $(W^{\prime},a)$ as in (2.8). If $\sigma$ is the Plücker permutations on $\{0,1,\dots,m\}$ corresponding to $\phi$ then $\phi b_{i}=b^{\sigma(i)}$. If $\mbox{WGr}^{\prime}(k,n)$ is the weighted Grassmann orbifold corresponding to $(W^{\prime},a)$ then $\mbox{WGr}^{\prime}(k,n)$ is homeomorphic to $\mbox{Gr}_{\sigma{\bf{b}}}(k,n)$. Thus Theorem 3.4 generalizes [7, Theorem 3.5].
Using Definition 2.3 it follows that if ${\bf b}=(b_{0},b_{1},\dots,b_{m})$ is a Plücker weight vector then $r{\bf b}:=(rb_{0},rb_{1},\dots,rb_{m})$ is also a Plücker weight vector for any positive integer $r$.
Proposition 3.6.
$\mbox{Gr}_{\bf b}(k,n)$ is equivariantly homeomorphic to $\mbox{Gr}_{r{\bf b}}(k,n)$ for every positive integer $r$.
Proof.
Consider the identity map ${\rm id}\colon Pl(k,n)\to Pl(k,n)$. The orbit map
$\pi_{\bf b}\colon Pl(k,n)\rightarrow\mbox{Gr}_{\bf b}(k,n)$
induces an (*$(\mathbb{C}^{*})^{n}$-equivariant homeomorphism $\mbox{Gr}_{\bf b}(k,n)\rightarrow\mbox{Gr}_{r{\bf b}}(k,n)$ for any positive integer $r$ such that the following diagram commutes. The (*$(\mathbb{C}^{*})^{n}$ action is defined in (3.1).
[…]
In [5], the authors have classified the weighted projective spaces $\mathbb{W}P(c_{0},\dots,c_{m})$ up to homeomorphism in terms of weight vectors.
A weight vector $(c_{0},c_{1},\dots,c_{m})\in(\mathbb{Z}_{\geq 0})^{m+1}$ is said to be a primitive weight vector if $\gcd\{c_{0},c_{1},\dots,c_{m}\}=1$. Moreover, A weight vector is called normalized if, for any prime $p$, at least two weights are not divisible by $p$. If a primitive weight vector is not normalized, then a prime $p$ exists such that all weights except one are divisible by $p$. Then, one can apply the normalization process by dividing all those vectors by $p$. One can apply the normalization process recurrently if the new vector is not normalized. Note that a repeated application normalization process can transform any weight vector into a unique normalized vector.
Proposition 3.7 ([5]).
Two weighted projective spaces $\mathbb{W}P(b_{0},\dots,b_{m})$ and $\mathbb{W}P(c_{0},\dots,c_{m})$ are homeomorphic iff $(b_{0},\dots,b_{m})$ and $(c_{0},\dots,c_{m})$ have same normalized form upto an order.
Remark 3.8.
Let ${\bf b}=(b_{0},b_{1},\dots,b_{m})$ and ${\bf c}=(c_{0},c_{1},\dots,c_{m})$ be two Plücker weight vectors such that $\mbox{Gr}_{\bf b}(k,n)$ and $\mbox{Gr}_{\bf c}(k,n)$ be homeomorphic. If we can extend this homeomorphism to get a homeomorphism between $\mathbb{W}P(b_{0},\dots,b_{m})$ and $\mathbb{W}P(c_{0},\dots,c_{m})$ then the normalized forms of $(b_{0},\dots,b_{m})$ and $(c_{0},\dots,c_{m})$ are the same up to an order.
In the remaining part of this section we study the rigidity problems of $\mbox{Gr}_{\bf b}(k,n)$ for $k\geq 2$. Define
$Ad(\lambda^{i}):=\{\lambda^{j}:|\lambda^{j}\cap\lambda^{i}|=k-1\}.$
Then cardinality of $Ad(\lambda^{i})$ is $k(n-k)$. Corresponding to $Ad(\lambda^{i})$ we define
${\bf b}_{(i)}:=\{{b_{j}}:j\in Ad(\lambda^{i})\}\text{ and }\overline{\bf b}_{(i)}:={\bf b}_{(i)}\cup\{b_{i}\}.$
Consider $[{e}_{\lambda^{i}}]\in\mbox{Gr}_{\bf b}(k,n)$, where $e_{\lambda^{i}}=(0,\dots,0,\underset{i^{th}}{1},0,\dots,0)\in Pl(k,n)$. We call the elements $[e_{\lambda^{i}}]$ by the coordinate elements, for $i=0,1,\dots,m$. Let
(3.4) $U_{i}=\{[z_{0},\dots,z_{m}]\in\mbox{Gr}_{\bf b}(k,n)~{}|~{}z_{i}\neq 0\}.$
From Section 2, it follows that $U_{i}\cong\mathbb{C}^{k(n-k)}/G(b_{i})$. Note that $U_{i}$ is an open cone over $L^{\prime}(b_{i};{\bf b}_{(i)})$. Let $t:=k(n-k)$. Using [13, Theorem 2] we have
$H^{q}(L^{\prime}(b_{i};{\bf{b}}_{(i)});\mathbb{Z})=\begin{cases}\mathbb{Z}&\quad\text{for}~{}q=0,2t-1\\ \mathbb{Z}_{\mu}&\quad\mu={l_{e}{\overline{\bf{b}}_{(i)}}}/{l_{e}{{\bf{b}}_{(i)}}}~{}\text{if}~{}q=2e~{}\text{for}~{}1\leq e\leq t-1\\ \{0\}&\quad\text{ otherwise}.\end{cases}$
For a prime $p$ let $r_{i}$ be the maximum non negative integer such that $p^{r_{i}}$ divides $b_{i}$ for all $i\in\{0,1,\dots,m\}$. We call $p^{r_{i}}$ as the $p$-th content in $b_{i}$. Let $p^{r_{1}},\dots,p^{r_{t}}$ be the $p$-th contents of the elements in ${\bf b}_{(i)}$ such that $r_{1}\leq\dots,\leq r_{t}$. Then
(3.5) $l_{e}{\bf{b}}_{(i)}=\prod_{p~{}\text{prime}}\prod_{i=1}^{e}p^{r_{t-i+1}}~{}\text{for}~{}1\leq e\leq t-1.$
Lemma 3.9.
If ${\bf b}$ is a primitive Plücker weight vector then $\overline{\bf{b}}_{(i)}$’s are primitive vectors for all $i=0,1,\dots,m$.
Proof.
Fix one $i\in\{0,1,\dots,m\}$. If $b_{j}\notin\overline{\bf b}_{(i)}$ then $|\lambda^{j}\cap\lambda^{i}|=k-u\leq k-2$. Without loss of generality assume that $\lambda^{i}=(\lambda^{i}_{1},\dots,\lambda^{i}_{u},\lambda^{i}_{u+1}\dots,\lambda^{i}_{k})$ and $\lambda^{j}=(\lambda^{j}_{1},\dots,\lambda^{j}_{u},\lambda^{i}_{u+1}\dots,\lambda^{i}_{k})$, where $\lambda^{i}_{1}\neq\lambda^{j}_{1},$ $\dots$,… , $\lambda^{i}_{u}\neq\lambda^{j}_{u}$. Consider Schubert symbols $\lambda^{s_{1}}=(\lambda^{j}_{1},\lambda^{i}_{2}\dots,\lambda^{i}_{u},\lambda^{i}_{u+1}\dots,\lambda^{i}_{k})$ $,\dots,\lambda^{s_{u}}=(\lambda^{i}_{1},\dots,\lambda^{i}_{u-1},\lambda^{j}_{u},\lambda^{i}_{u+1},\dots,\lambda^{i}_{k})$. Then $|\lambda^{i}\cap\lambda^{s_{v}}|=k-1$ and $b_{s_{v}}\in{\bf b}_{(i)}$ for all $v=1,2,\dots,u$. Note that
$\bigcup_{v=1}^{u}\lambda^{s_{v}}=\lambda^{j}\bigcup_{(u-1)\text{ times}}\lambda^{i}.$
Since ${\bf b}$ is a Plücker weight vector using remark 2.4, there exist $b_{s_{1}},b_{s_{2}},\dots,b_{s_{u}}\in{\bf b}_{(i)}$ such that
$b_{j}+(u-1)b_{i}=b_{s_{1}}+b_{s_{2}}\dots+b_{s_{u}}.$
If $d$ divides $b_{i}$ and every element of ${\bf b}_{(i)}$ then $d$ also divides $b_{j}$. Thus $\overline{\bf b}_{(i)}$ is not a primitive vector contradicts that ${\bf b}$ is a primitive Plücker weight vector. ∎
Lemma 3.10.
If ${\bf b}$ is a primitive Plücker weight vector then
$H^{2t-1}({\rm{Gr}}_{\bf b}(k,n),{\rm{Gr}}_{\bf b}(k,n)\setminus[e_{\lambda^{i}}]~{};~{}\mathbb{Z})=\mathbb{Z}_{b_{i}}.$
Proof.
Recall $U_{i}$ from (3.4). Using excision
$H^{2t-1}(\mbox{Gr}_{\bf b}(k,n),\mbox{Gr}_{\bf b}(k,n)\setminus[e_{\lambda^{i}}]~{};~{}\mathbb{Z})\cong H^{2t-1}(U_{i},U_{i}\setminus[e_{\lambda^{i}}]~{};~{}\mathbb{Z}).$
Now using the fact that $U_{i}$ is an open cone over $L^{\prime}(b_{i};{\bf b}_{(i)})$
$H^{2t-1}(U_{i},U_{i}\setminus[e_{\lambda^{i}}]~{};~{}\mathbb{Z})\cong H^{2t-2}(L(b_{i};{\bf{b}}_{(i)});\mathbb{Z}).$
Now we need to prove
(3.6) $\frac{l_{t-1}\overline{\bf{b}}_{(i)}}{l_{t-1}{\bf{b}}_{(i)}}=b_{i}.$
Since ${\bf b}$ is primitive, we have $\overline{\bf b}_{(i)}$ is also primitive for all $i=0,1,\dots,m$, using Lemma 3.9. For a prime $p$, if $p$ does not divide $b_{i}$, then the $p$-th content of both sides of the (3.6) is 1. Now let $p$ divides $b_{i}$. Since $\overline{\bf b}_{(i)}$ is primitive there exists at least one element $b_{j}$ in ${\bf b}_{(i)}$ such that $p$ does not divide $b_{j}$. Note that for every element $b_{j}$ in ${\bf b}_{(i)}$ there exist $b_{j^{\prime}},b_{s},b_{s^{\prime}}\in{\bf b}_{(i)}$ such that
$b_{j}+b_{j^{\prime}}=b_{s}+b_{s^{\prime}}.$
Thus there exist at least two integers in ${\bf b}_{(i)}$ such that both integers are not divisible by $p$. Then $p$-th content of $b_{i}$=$p$-th content of $\frac{l_{t-1}\overline{\bf{b}}_{(i)}}{l_{t-1}{\bf{b}}_{(i)}}$. This is true for all prime $p$. This completes the proof. ∎
Theorem 3.11.
If there exists a homeomorphism $f$ between two weighted Grassmann orbifolds ${\rm{Gr}}_{\bf b}(k,n)$ and ${\rm{Gr}}_{\bf c}(k,n)$ such that for all $i=0,1,\dots,m$, $f$ maps every coordinate element $[e_{\lambda^{i}}]$ to some coordinate element $[e_{\lambda^{j}}]$ then ${\bf b}$ and ${\bf c}$ are same up to a scalar multiplication and a permutation.
Proof.
Without loss of generality, we can assume that both the Plücker weight vectors ${\bf b}$ and ${\bf c}$ are primitive. Let $f$ be the homeomorphism between $\mbox{Gr}_{\bf b}(k,n)$ and $\mbox{Gr}_{\bf c}(k,n)$ such that for all $i\in\{0,1,\dots,m\}$, $f([e_{\lambda^{i}}])=[e_{\lambda^{j}}]$ for some $j\in\{0,1,\dots,m\}$. Then it induces a isomorphism between $H^{2t-1}(\mbox{Gr}_{\bf b}(k,n),\mbox{Gr}_{\bf b}(k,n)\setminus[e_{\lambda^{i}}];\mathbb{Z})$ and $H^{2t-1}(\mbox{Gr}_{\bf c}(k,n),\mbox{Gr}_{\bf c}(k,n)\setminus[e_{\lambda^{j}}];\mathbb{Z})$. This implies $\mathbb{Z}_{b_{i}}$ is isomorphic to $\mathbb{Z}_{c_{j}}$. Thus $b_{i}=c_{j}$. Define a permutation $\sigma$ on $\{0,1,\dots,m\}$ by $\sigma(i)=j$. Then ${\bf b}=\sigma{\bf c}$ for some permutation $\sigma$ on $\{0,1,\dots,m\}$. ∎
We recall the (*$(\mathbb{C}^{*})^{n}$-action on $\mbox{Gr}_{\bf b}(k,n)$ from (3.1) and note that the fixed point set of $\mbox{Gr}_{\bf b}(k,n)$ with respect to the (*$(\mathbb{C}^{*})^{n}$-action is $\{[e_{\lambda^{i}}]\}_{i=0}^{m}$.
Corollary 3.12.
[…]
Proof.
A weakly equivariant homeomorphism maps a fixed point to a fixed point. The coordinate elements $\{[e_{\lambda^{i}}]\}_{i=0}^{m}$ are the fixed point set in $\mbox{Gr}_{\bf b}(k,n)$ with respect to the action of $G$. Hence, the proof follows from the Theorem 3.11. ∎
4. Torsion’s in the integral cohomology of weighted Grassmann orbifold
In this section, we prove that for a prime $p$ the integral cohomologies of weighted Grassmann orbifolds have no $p$-torsion under some assumptions. We prove that for any Plücker weight vector ${\bf b}\in(\mathbb{Z}_{\geq 1})^{6}$, $H_{i}(G_{\bf b}(k,n))$ is torsion free if $i\neq 3$. We discuss some non-trivial classes of weighted Grassmann orbifolds whose integral cohomologies have no torsions. In particular, the integral cohomology of a divisive weighted Grassmann orbifold is torsion-free and concentrated in even degrees.
Theorem 4.1.
For a prime $p$, if there exists a Plücker permutation $\sigma$ on $\{0,\dots,m\}$ such that the for all $j\geq 3$, the $p$-th content of $b_{\sigma(j)}$ divides at least $d(\lambda^{j})-1$ many elements of the set $\{b_{\sigma(\ell)}:\lambda^{\ell}\in R(\lambda^{j})\}$ then the integral cohomology of ${\rm{Gr}}_{\bf b}(k,n)$ has no $p$-torsion.
Proof.
Consider the Plücker permutation $\sigma$ and recall the building sequence
$\sigma X^{0}\subset\sigma X^{1}\subset\dots\subset\sigma X^{m}=\mbox{Gr}_{\bf b}(k,n),$
induced from the $q$-CW complex structure as in Theorem 3.4, where $X^{i}$ is defined in (2.12). Note that $\sigma X^{i}$ is weighted projective space for all $0\leq i\leq 2$. Thus, the integral cohomology of $\sigma X^{i}$ has no torsion for all $0\leq i\leq 2$ by [13, Theorem 1]. We complete the proof by induction on $i$. Let $H^{*}(\sigma X^{i-1};\mathbb{Z})$ has no $p$-torsion. Consider the cofibration as in Proposition 2.12
$L^{\prime}(b_{\sigma(i)};\sigma{\bf{b}}^{(i)})\hookrightarrow\sigma X^{i-1}\to\sigma X^{i},$
where ${\bf b}^{(i)}$ is defined in (2.13). This will induce the following exact sequence in cohomology
(4.1) $\to\widetilde{H}^{q-1}(L^{\prime}(b_{\sigma(i)};\sigma{\bf{b}}^{(i)});\mathbb{Z})\to H^{q}(\sigma X^{i};\mathbb{Z})\to H^{q}(\sigma X^{i-1};\mathbb{Z})\to\widetilde{H}^{q}(L^{\prime}(b_{\sigma(i)};\sigma{\bf{b}}^{(i)});\mathbb{Z}).$
Using [13, Theorem 2] we have
$H^{q}(L^{\prime}(b_{\sigma(i)};\sigma{\bf{b}}^{(i)});\mathbb{Z})=\begin{cases}\mathbb{Z}&\quad\text{for}~{}q=0,2d(\lambda^{i})-1\\ \mathbb{Z}_{\mu}&\quad\mu={l_{e}{\sigma\overline{\bf{b}}^{(i)}}}/{l_{e}{\sigma{\bf{b}}^{(i)}}}~{}\text{if}~{}q=2e~{}\text{for}~{}1\leq e\leq d(\lambda^{i})-1\\ \{0\}&\quad\text{ otherwise},\end{cases}$
where ${\overline{\bf{b}}^{(i)}}:={\bf{b}}^{(i)}\cup\{b_{i}\}.$
For any prime $p$ the $p$-th content in $l_{e}{\sigma{\bf{b}}^{(i)}}$ is the product of $e$ many maximum $p$-th contents of the elements in set $\{b_{\sigma(\ell)}\colon\lambda^{\ell}\in R(\lambda^{i})\}$ as in (3.5). Thus the $p$-th content of $\mu=1$ if the $p$-th content of $b_{\sigma(i)}$ divides $e$ many element of the set $\{b_{\sigma(\ell)}:\lambda^{\ell}\in R(\lambda^{i})\}$. Now the maximum value of $e$ can be $d(\lambda^{i})-1$. Thus by our assumption $H^{*}(L^{\prime}(b_{\sigma(i)};\sigma{\bf{b}}^{(i)});\mathbb{Z})$ has no $p$-torsion. Thus $H^{*}(\sigma X^{i},\mathbb{Z})$ has no $p$-torsion using (4.1). Hence, by induction the integral cohomology of $\mbox{Gr}_{\bf b}(k,n)$ has no $p$-torsion. ∎
Remark 4.2.
If Theorem 4.1 is true for every prime $p$ then $H^{*}(\mbox{Gr}_{\bf{b}}(k,n),\mathbb{Z})$ is torsion-free and contrated in even degrees.
Recall the subcomplexes $X^{i}$ of $\mbox{Gr}_{\bf{b}}(k,n)$ from (2.12) for $i=0,1,\dots,m$. Define
$\eta=\frac{\operatorname{lcm}\{b_{0},b_{1},b_{3}\}}{\operatorname{lcm}\{b_{0},b_{1}\}}\text{ and }\eta^{\prime}=\frac{\operatorname{lcm}\{b_{0},b_{1},b_{2}\}}{\operatorname{lcm}\{b_{0},b_{1}\}}.$
Lemma 4.3.
$H_{i}(X^{3};\mathbb{Z})$ has no torsion if $i\neq 2$. Moreover if $\gcd\{\eta,\eta^{\prime}\}=1$ then $H_{i}(X^{3};\mathbb{Z})$ is torsion-free for all $i$.
Proof.
Consider the cofibration
$L^{\prime}(b_{3};(b_{0},b_{1}))\to\mathbb{W}P(b_{0},b_{1},b_{2})\to X^{3}.$
This induces an exact sequence in homology by the following
(4.2) $\to\widetilde{H}_{i}(L^{\prime}(b_{3};(b_{0},b_{1})))\to H_{i}(\mathbb{W}P(b_{0},b_{1},b_{2}))\to H_{i}(X^{3})\to\widetilde{H}_{i-1}(L^{\prime}(b_{3};(b_{0},b_{1})))\to$
For $i=1$ in (4.2) we get
$0\to H_{1}(X^{3})\to\mathbb{Z}\to\mathbb{Z}\to.$
$H_{1}(X^{3})$ is either $\{0\}$ or a finite abelian group as $X^{3}$ has no cell of real dimension $1$. Also $H_{1}(X^{3})\to\mathbb{Z}$ is injective. Thus $H_{1}(X^{3})=\{0\}$.
For $i=3$ in (4.2) we get
$\to\mathbb{Z}\to 0\to H_{3}(X^{3})\to 0\to\mathbb{Z}.$
Thus $H_{3}(X^{3})=\{0\}$. For $i=4$ in (4.2) we get
$\to 0\to\mathbb{Z}\to H_{4}(X^{3})\to\mathbb{Z}\to 0.$
Thus $H_{4}(X^{3})=\mathbb{Z}\oplus\mathbb{Z}$. For $i\geq 5$ in (4.2) we get
$\to 0\to 0\to H_{i}(X^{3})\to 0\to.$
Now put $i=2$ in (4.2) we get
$\to 0\to\mathbb{Z}\to H_{2}(X^{3})\to\mathbb{Z}_{\eta}\to 0,$
Thus, torsion can arise only in $H_{2}(X^{3})$, and there can be maximum $\eta$ torsion. There exist another cofibration
$L^{\prime}(b_{2};(b_{0},b_{1}))\to\mathbb{W}P(b_{0},b_{1},b_{3})\to X^{3}.$
Considering the homology exact sequence corresponding to this cofibration, we can say $H_{2}(X^{3})$ has maximum $\eta^{\prime}$ torsion. Hence If $\gcd(\eta,\eta^{\prime})=1$ then $H_{i}(X^{3},\mathbb{Z})$ has no torsion for all $i$. ∎
Corollary 4.4.
For a prime $p$, there always exist a Plücker permutation $\sigma$ on $\{0,1,\dots,m\}$ such that $H_{i}(\sigma X^{3};\mathbb{Z})$ has no $p$-torsion for all $i$.
Theorem 4.5.
$H^{i}({\rm{Gr}}_{\bf{b}}(2,4);\mathbb{Z})$ is torsion free for any weighted Grassmann orbifold ${\rm{Gr}}_{\bf b}(2,4)$ if $i\neq 3$.
Proof.
For a prime $p$, choose a Plücker permutation $\sigma$ such that the minimum power of $p$ attains at $b_{\sigma(5)}$. Note that there are $4$ different Plücker permutation which satisfy the above condition. We can choose $\sigma$ such that $r^{\sigma(4)}\leq\{r^{\sigma(1)},r^{\sigma(2)},r^{\sigma(3)},r^{\sigma(4)}\}$. From the cofibration $L^{{}^{\prime}}(b_{\sigma(5)};\sigma{\bf{b}}^{(5)})\to\sigma X^{4}\hookrightarrow\mbox{Gr}_{\bf{b}}(2,4)$ we have
$\to H_{i}(\sigma X^{4})\to H_{i}(\mbox{Gr}_{\bf{b}}(2,4))\to\widetilde{H}_{i-1}(L^{\prime}(b_{\sigma(5)};\sigma{\bf{b}}^{(5)}))\to H_{i-1}(\sigma X^{4})\to.$
For $i=8$, we have
$\to 0\to H_{8}(\mbox{Gr}_{\bf{b}}(2,4))\to\mathbb{Z}\to 0\to.$
For $2\leq i\leq 7$, $\widetilde{H}_{i-1}(L^{\prime}(b_{\sigma(5)};\sigma{\bf{b}}^{(5)}))$ has no $p$-torsion. Thus $H_{i}(\mbox{Gr}_{\bf{b}}(2,4))$ has no $p$-torsion if the same is true for $H_{i}(\sigma X^{4})$. Similarly from the cofibration $L^{{}^{\prime}}(b_{\sigma(4)},\sigma{\bf{b}}^{(4)})\to\sigma X^{3}\hookrightarrow\sigma X^{4}$ we have
$\to\widetilde{H}_{i}(L^{\prime}(b_{\sigma(4)};\sigma{\bf{b}}^{4}))\to H_{i}(\sigma X^{3})\to H_{i}(\sigma X^{4})\to\widetilde{H}_{i-1}(L^{\prime}(b_{\sigma(4)};\sigma{\bf{b}}^{4}))\to.$
For $i=6$ we
$0\to H_{6}(\sigma X^{4})\to\mathbb{Z}\to 0$
Thus $H_{6}(\sigma X^{4})=\mathbb{Z}$. For $2\leq i\leq 5$, $H_{i}(\sigma X^{4})$ has no $p$-torsion if the same is true for $H_{i}(\sigma X^{3})$ as $\widetilde{H}_{i-1}(L^{\prime}(b_{\sigma(4)};\sigma{\bf{b}}^{(4)}))$ has no $p$-torsion. Using Lemma 4.3, $H_{i}(\sigma X^{3})$ has no torsion if $i\neq 2$. Thus $H_{i}(\mbox{Gr}_{\bf{b}}(2,4);\mathbb{Z})$ has no torsion if $i\neq 2$. Hence the result follows using the universal coefficient Theorem for cohomology. ∎
Remark 4.6.
If $H_{2}(\sigma X^{3};\mathbb{Z})$ has no $p$-torsion for the same $\sigma$ that requires in the proof of Theorem 4.5 then $H_{i}(\mbox{Gr}_{\bf{b}}(2,4);\mathbb{Z})$ has no $p$-torsion for every $i$.
Now we show that the theorem 4.1 is applicable to a huge class of weighted Grassmann orbifolds.
Example 4.7.
Let $n=4,k=2$ and $m={4\choose 2}-1=5$. In this case, we have $6$ Schubert symbols given by
$\lambda^{0}=(1,2)<\lambda^{1}=(1,3)<\lambda^{2}=(1,4)<\lambda^{3}=(2,3)<\lambda^{4}=(2,4)<\lambda^{5}=(3,4).$
  1. (1)
    Let $\alpha,\beta,\gamma$ be three positive integers. Consider
    ${\bf{b}}=(\alpha\beta\gamma,\alpha\beta\gamma,(\alpha\beta-\alpha+1)\gamma,\alpha\gamma,\gamma,\gamma)\in(\mathbb{Z}_{\geq 1})^{6}.$
    Then $b_{i}$ divides $b_{j}$ for all $\lambda^{j}\preceq\lambda^{i}$ for $i\geq 3$. Thus the integral cohomology of $\mbox{Gr}_{\bf{b}}(2,4)$ has no torsion using Theorem 4.1.
  2. (2)
    Let $\alpha,\beta,\gamma,s$ and $t$ be positive integers. Consider a weight vector of the form
    ${\bf{b}}=(\gamma,(1+\beta)st\alpha,(1+s\beta)t\alpha,st\alpha,t\alpha,\alpha)\in(\mathbb{Z}_{\geq 1})^{6}.$
    In addition, if $\alpha+\gamma=t\alpha(1+s+s\beta)$ then ${\bf{b}}$ is a Plücker weight vector. The Plücker weight vector ${\bf{b}}$ satisfies the assumption of Theorem 4.1. Thus the integral cohomology of the corresponding $\mbox{Gr}_{\bf{b}}(2,4)$ has no torsion using Theorem 4.1.
Theorem 4.8.
[…]
Proof.
Since ${\bf{b}}$ is a Plücker weight vector we have $b_{1}+b_{4}=b_{2}+b_{3}$. If $b_{1}=b_{2}$ then $b_{3}=b_{4}$. Thus ${\bf{b}}=(b_{0},b_{1},b_{1},b_{3},b_{3},b_{5})$. For a prime $p$ we can assume $p$-th content of $b_{3}$ divides $p$-th content of $b_{1}$. Then by Theorem 4.1, $H^{*}(X^{4};\mathbb{Z})$ has no torsion. Otherwise consider the Plücker permutation $\sigma$ defined by $\sigma(1)=4$, $\sigma(4)=1$ and $\sigma(i)=i$ for $i\neq 1,4$. Then $\sigma{\bf{b}}=(b_{0},b_{3},b_{3},b_{1},b_{1},b_{5})$. Consider the weighted Grassmann orbifold $\mbox{Gr}_{\sigma{\bf{b}}}(2,4)$ to ensure that $H^{*}(X^{4};\mathbb{Z})$ has no torsion. Using the cohomology exact sequence in (4.1), $H^{*}(\mbox{Gr}_{\bf{b}}(2,4);\mathbb{Z})$ has maximum $b_{5}$ torsion. Consider the Plücker permutation $\sigma^{\prime}$ defined by $\sigma^{\prime}(0)=5$, $\sigma^{\prime}(5)=0$ and $\sigma^{\prime}(i)=i$ for $i\neq 0,5$. Then $\sigma^{\prime}{\bf{b}}=(b_{5},b_{1},b_{1},b_{3},b_{3},b_{0})$. By a similar argument $H^{*}(\mbox{Gr}_{\sigma^{\prime}{\bf{b}}}(2,4);\mathbb{Z})$ has maximum $b_{0}$ torsion. Since $(b_{0},b_{5})=1$ then $H^{*}(\mbox{Gr}_{\bf{b}}(2,4);\mathbb{Z})$ has no torsion. ∎
In the following example we discuss some classes of $\mbox{Gr}_{\bf{b}}(2,5)$ such that $H^{*}(\mbox{Gr}_{\bf{b}}(2,5);\mathbb{Z})$ has no torsion.
Example 4.9.
Let $n=5,k=2$ and $m={5\choose 2}-1=9$. In this case, we have $10$ Schubert symbols given by
$\lambda^{0}=(1,2)<\lambda^{1}=(1,3)<\lambda^{2}=(1,4)<\lambda^{3}=(1,5)<\lambda^{4}=(2,3)<$
$\lambda^{5}=(2,4)<\lambda^{6}=(2,5)<\lambda^{7}=(3,4)<\lambda^{8}=(3,5)<\lambda^{9}=(4,5).$
  1. (1)
    Let $\alpha,\beta,\gamma$ be three positive integers. Consider
    ${\bf{b}}=(\alpha\beta\gamma,\alpha\beta\gamma,(\alpha\beta-\alpha+1)\gamma,\alpha\beta\gamma,\alpha\gamma,\gamma,\alpha\gamma,\gamma,\alpha\gamma,\gamma)\in(\mathbb{Z}_{\geq 1})^{10}.$
    Then the integral cohomology of $\mbox{Gr}_{\bf{b}}(2,5)$ has no torsion using Theorem 4.1.
  2. (2)
    Let $\alpha,\beta,\gamma,s$ ant $t$ be positive integers. Consider
    ${\bf{b}}=(\gamma,(1+\beta)st\alpha,(1+s\beta)t\alpha,(1+\beta)st\alpha,st\alpha,t\alpha,t\alpha,\alpha,\alpha,\alpha)\in(\mathbb{Z}_{\geq 1})^{10}.$
    Then the integral cohomology of $\mbox{Gr}_{\bf{b}}(2,5)$ has no torsion using Theorem 4.1.
  3. (3)
    Let $a$ and $c$ be two positive integer such that $\gcd\{a,c\}=1$. Let $s=a+c$. For any positive integer $\alpha$ such that $s-\alpha$ divides $\alpha$ and $c$ divides $\alpha$ then the $\mbox{Gr}_{\bf{b}}(2,5)$ has no torsion where ${\bf{b}}=(a,\alpha,\alpha,\alpha,s-\alpha,s-\alpha,s-\alpha,c,c,c)$. For example, ${\bf{b}}=(5,4,4,4,2,2,2,1,1,1)$, ${\bf{b}}=(28,24,24,24,6,6,6,2,2,2)$
Now we define divisive weighted Grassmann orbifold in terms of Plücker weight vector and Plücker permutation following [7].
Definition 4.10.
A weighted Grassmann orbifold $\mbox{Gr}_{\bf b}(k,n)$ is called divisive if there exists a Plücker permutation $\sigma$ on $\{0,1,\dots,m\}$ such that $b_{\sigma(i)}$ divides $b_{\sigma(i-1)}$ for every $i=1,2,\dots,m$, where $m+1={n\choose k}$.
Example 4.11.
Let $\alpha$ and $\beta$ be two positive integers. Consider a Plücker weight vector ${\bf b}=(\alpha,\alpha\beta,\alpha\beta,\alpha,\alpha,\alpha\beta)$. Define a Plücker permutation $\sigma$ on $\{0,1,\dots,5\}$ by $\sigma(0)=5$, $\sigma(5)=0$ and $\sigma(i)=i$ for $i=1,2,3,4$. Then $\mbox{Gr}_{\bf b}(2,4)$ is a divisive weighted Grassmann orbifold.
The proof of the following theorem follows from [7, Theorem 3.19].
Theorem 4.12.
A divisive weighted Grassmann orbifold has a CW complex structure with even dimensional cells $\{\mathbb{C}^{d(\lambda^{i})}\}_{i=0}^{m}$.
Corollary 4.13.
Integral cohomologies of divisive weighted Grassmann orbifolds are torsion-free and concentrated in even degrees.
5. Structure constants for the equivariant cohomology rings of divisive weighted Grassmann orbifolds with integer coefficients
In this section, first we revisit the $T^{n}$-equivariant cohomology ring of divisive weighted Grassmann orbifolds with integer coefficients as introduced in [7]. Our focus lies on studying the multiplication formulae in the integral equivariant cohomology ring of a divisive weighted Grassmann orbifold. We consider a basis of the aforementioned ring as a $\mathbb{Z}[y_{1},\dots,y_{n}]$ module and compute all the equivariant structure constants concerning that basis. Moreover, we demonstrate the positivity of the equivariant structur constants.
Let $T^{n}=(S^{1})^{n}$ be the $n$-dimensional compact abelian torus. Consider the $T^{n}$-action on $\mbox{Gr}_{\bf{b}}(k,n)$ as the restriction of the (*$(\mathbb{C}^{*})^{n}$-action defined in (3.1). Let $\mbox{Gr}_{\bf b}(k,n)$ be a divisive weighted Grassmann orbifold. We can assume that $b_{i}$ divides $b_{i-1}$ for all $i=1,2,\dots,m$ using Theorem 3.4. Define $d_{ij}:=\frac{b_{i}}{b_{j}}$ for $i<j$ and recall $R(\lambda^{i})$ from (2.11).
Theorem 5.1.
[7, Theorem 4.7] The $T^{n}$-equivariant cohomology ring of a divisive weighted Grassmann orbifold ${\rm{Gr}}_{\bf{b}}(k,n)$ with integer coefficients is given by
[…]
Here $\mathbb{Z}[y_{1},\dots,y_{n}]=H^{*}_{T^{n}}(\{pt\};\mathbb{Z})$ and $Y_{\lambda}:=\sum_{i=1}^{k}y_{\lambda_{i}}$ for the Schubert symbol $\lambda=(\lambda_{1},\ldots,\lambda_{k})$.
For the particular case, ${\bf{b}}=(1,1,\dots,1)$ the $T^{n}$-equivariant cohomology ring of $\mbox{Gr}_{\bf b}(k,n)$ coincide with the $T^{n}$-equivariant cohomology ring of $\mbox{Gr}(k,n)$ considered in [14]. In [14, Section 2], the authors computed the multiplication formulae for the integral equivariant cohomology ring $H_{T^{n}}^{*}(\mbox{Gr}(k,n);\mathbb{Z})$ of the Grassmann manifold $\mbox{Gr}(k,n)$.
5.1. Knutson-Tao’s results on $H_{T^{n}}^{*}(\mbox{Gr}(k,n);\mathbb{Z})$
Corresponding to every Schubert symbol $\lambda=(\lambda_{1},\dots,\lambda_{k})$, one can consider a string consisting of $k$ ones and $(n-k)$ zeroes in the following order. One for each $\lambda_{i}$-th position and zero for the remaining $(n-k)$ positions. In this way, we can identify the set of all Schubert symbols $\lambda=(\lambda_{1},\dots,\lambda_{k})$ and the set of strings consisting of $k$ ones and $(n-k)$ zeroes in arbitrary order, as explored by Knutson and Tao, see [14]. We recall an inversion of a Schubert symbol $\lambda$ is a pair $(s,s^{\prime})$ such that $s\in\lambda,s^{\prime}\notin\lambda$ and $s<s^{\prime}$. We denote the set of all inversions of $\lambda$ by $\rm{inv}(\lambda)$. If $(s,s^{\prime})\in\rm{inv}(\lambda)$, then $(s,s^{\prime})\lambda$ denote the Schubert symbol obtained by replacing $s$ by $s^{\prime}$ in $\lambda$ and ordering the later set.
For every $i\in\{0,1,\dots,m\}$ consider a subset of Schubert symbols
$I(\lambda^{i})=\{\lambda^{j}:\lambda^{j}=(s,s^{\prime})\lambda^{i}\text{ for }(s,s^{\prime})\in\rm{inv}(\lambda^{i})\}.$
Let $d^{\prime}(\lambda^{i})$ denotes the cardinality of the set $I(\lambda^{i})$. Then $d(\lambda^{i})+d^{\prime}(\lambda^{i})=k(n-k)$.
Proposition 5.2.
[14] There exist an equivariant Schubert basis $\{\widetilde{S}_{\lambda^{i}}\}_{i=0}^{m}$ for the equivariant cohomology ring $H_{T^{n}}^{*}(\mbox{Gr}(k,n);\mathbb{Z})$. Each basis element $\widetilde{S}_{\lambda^{i}}$ is uniquely determined the following conditions
  1. (1)
    $\widetilde{S}_{\lambda^{i}}|_{\lambda^{j}}\neq 0\implies\lambda^{j}\preceq\lambda^{i}$.
  2. (2)
    $\widetilde{S}_{\lambda^{i}}|_{\lambda^{i}}=\prod_{\lambda^{\ell}\in I(\lambda^{i})}(Y_{\lambda^{\ell}}-Y_{\lambda^{i}})$.
  3. (3)
    $\widetilde{S}_{\lambda^{i}}|_{\lambda^{j}}$ is a homogeneous polynomial of $y_{1},y_{2},\dots,y_{n}$ of degree $d^{\prime}({\lambda^{i}})$.
We have the following multiplication rule in the equivariant cohomology ring $H_{T^{n}}^{*}(\mbox{Gr}(k,n);\mathbb{Z})$
(5.1) $\widetilde{S}_{\lambda^{i}}\widetilde{S}_{\lambda^{j}}=\sum_{\lambda^{\ell}}\widetilde{C}_{i~{}j}^{\ell}\widetilde{S}_{\lambda^{\ell}}.$
In their study [14], Knutson and Tao conducted explicit computations of the equivariant structure constant $\widetilde{C}_{i~{}j}^{\ell}$ with respect to the basis $\{\widetilde{S}_{\lambda^{i}}\}$ for $H_{T^{n}}^{*}(\mbox{Gr}(k,n);\mathbb{Z})$. We refer to [14, Section 1] for the definition of puzzle pieces and equivariant puzzle pieces made by triangles and rhombus with all edges labeled 0 or 1. One can use these pieces to construct a puzzle $P$ as a decomposition of a labeled equilateral triangle into puzzle pieces and equivariant puzzle pieces. By a labeled equilateral triangle, we mean an upward-pointing equilateral triangle with integer side length $n$ and each side is labeled by a string consisting of $k$ ones and $(n-k)$ zeroes. The labeled equilateral triangle induced from the puzzle $P$ is denoted by $\partial P$, the boundary of $P$. Let $P$ be a puzzle with $\partial P=\Delta_{\lambda^{i}~{}\lambda^{j}}^{\lambda^{\ell}}$. Using [14] the number of equivariant puzzle pieces in $P$ is $|P|:=d^{\prime}(\lambda^{i})+d^{\prime}(\lambda^{j})-d^{\prime}(\lambda^{\ell})\geq 0$. Let $p_{s}$ be equivariant puzzle pieces of $P$, for $s=1,2,\dots,|P|.$ Then $wt(P)\colon=\prod_{s=1}^{|P|}wt(p_{s})$, where $wt(p_{s})=y_{a}-y_{c}$ for $a>c$ and $wt(P)=1$ if $|P|=0$.
Proposition 5.3.
[14, Theorem 2] The equivariant structure constant is equal to the sum of the weights of all puzzle $P$ such that $\partial P=\Delta_{\lambda^{i}~{}\lambda^{j}}^{\lambda^{\ell}}$.
$\widetilde{C}_{i~{}j}^{\ell}=\sum_{P~{}:~{}\partial P=\Delta_{\lambda^{i}~{}\lambda^{j}}^{\lambda^{\ell}}}wt(P)$
5.2. Structure constants for the equivariant cohomology of $\mbox{Gr}_{\bf b}(k,n)$
In this subsection, we compute the multiplication structure in the equivariant cohomology ring $H_{T^{n}}^{*}(\mbox{Gr}_{\bf b}(k,n);\mathbb{Z})$.
Proposition 5.4.
[7] There exist an equivariant Schubert basis $\{{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{m}$ for the equivariant cohomology ring $H_{T^{n}}^{*}(\mbox{Gr}_{\bf{b}}(k,n);\mathbb{Z})$ of a divisive weighted Grassmann orbifold. Moreover, each basis element ${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}$ is uniquely determined by the following conditions.
  1. (1)
    ${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}|_{\lambda^{j}}\neq 0\implies\lambda^{j}\succeq\lambda^{i}$.
  2. (2)
    ${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}|_{\lambda^{i}}=\prod_{\lambda^{\ell}\in R(\lambda^{i})}(Y_{\lambda^{\ell}}-\dfrac{b_{\ell}}{b_{i}}Y_{\lambda^{i}})$.
  3. (3)
    ${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}|_{\lambda^{j}}$ is a homogeneous polynomial of $y_{1},y_{2},\dots,y_{n}$ of degree $d({\lambda^{i}})$.
Subsequently, we have the following product formulae in the equivariant cohomology ring of divisive weighted Grassmann orbifold with integer coefficients.
(5.2) ${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{j}}=\sum_{\ell=0}^{m}{\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{\ell}{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{\ell}}$
Our goal is to compute the equivariant structure constant ${\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{\ell}$ explicitly. The author and Sarkar have studied some interesting property of ${\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{\ell}$ in [7, Lemma 5.5].
For the equivariant cohomology ring $H^{*}_{T^{n}}(\mbox{Gr}(k,n);\mathbb{Z})$ we denote the basis $\{{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{m}$ by $\{\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{m}$. Let $\sigma_{r}\in S_{n}$ be the permutation defined by
$\sigma_{r}:=\begin{pmatrix}1&2&3&\cdots&n-1&n\\ n&n-1&n-2&\cdots&2&1\end{pmatrix}.$
In other words, $\sigma_{r}(i)=n+1-i$. Note that $\sigma_{r}^{2}(i)=i$. Then $\sigma_{r}$ induces a permutation on $\mathbb{Z}[y_{1},y_{2},\dots,y_{n}]$ defined by $\sigma_{r}y_{i}=y_{n+1-i}$. Thus $\sigma_{r}$ induces a permutation on $H_{T^{n}}(\mbox{Gr}_{\bf{b}}(k,n);\mathbb{Z})$.
The following proposition gives the relation between two bases of $H_{T^{n}}(\mbox{Gr}_{\bf{b}}(k,n);\mathbb{Z})$.
Proposition 5.5.
$\widetilde{\mathbb{S}}_{\sigma_{r}\lambda^{i}}|_{\sigma_{r}\lambda^{j}}=\sigma_{r}(\widetilde{S}_{\lambda^{i}}|_{\lambda^{j}}).$
Proof.
Note that $\lambda^{\ell}\in R(\sigma_{r}\lambda^{i})$ if and only if $\sigma_{r}\lambda^{\ell}\in I(\lambda^{i})$. Thus
$\widetilde{\mathbb{S}}_{\sigma_{r}\lambda^{i}}|_{\sigma_{r}\lambda^{i}}$ $=\prod_{\lambda^{\ell}\in R(\sigma_{r}\lambda^{i})}(Y_{\lambda^{\ell}}-Y_{\sigma_{r}\lambda^{i}})$
$=\prod_{\sigma_{r}\lambda^{\ell}\in I(\lambda^{i})}(Y_{\sigma_{r}(\sigma_{r}\lambda^{\ell})}-Y_{\sigma_{r}\lambda^{i}})$
$=\prod_{\lambda^{s}\in I(\lambda^{i})}(Y_{\sigma_{r}\lambda^{s}}-Y_{\sigma_{r}\lambda^{i}})=\sigma_{r}(\widetilde{S}_{\lambda^{i}}|_{\lambda^{i}}).$
Also, both are of degree $d(\sigma_{r}\lambda^{i})=d^{\prime}(\lambda^{i})$. Now the proof follows from the uniqueness of the equivariant Schubert basis $\widetilde{\mathbb{S}}_{\lambda^{i}}$. ∎
Remark 5.6.
$\widetilde{S}_{\lambda^{i}}$ represents the equivariant cohomology class in $H_{T^{n}}^{*}(\mbox{Gr}(k,n);\mathbb{Z})$ corresponding to the closure of the cell $E(\lambda^{i})$, see [14, Section 2.3]. Using Proposition 5.5, $\widetilde{\mathbb{S}}_{\lambda^{i}}|_{\lambda^{j}}=\sigma_{r}(\widetilde{S}_{\sigma_{r}\lambda^{i}}|_{\sigma_{r}\lambda^{j}}).$ Thus one can think $\widetilde{\mathbb{S}}_{\lambda^{i}}$ as the equivariant cohomology class in $H_{T^{n}}^{*}(\mbox{Gr}(k,n);\mathbb{Z})$ corresponding to the closure of the cell $\sigma_{r}E(\sigma_{r}\lambda^{i})$.
Let $\widetilde{\mathscr{C}}_{i~{}j}^{\ell}$ be the equivariant structure constant corresponding to the basis $\{\widetilde{\mathbb{S}}_{\lambda^{i}}\}$. Then we have the following formulae for $\widetilde{\mathscr{C}}_{i~{}j}^{\ell}$. Let $p_{s}$ be an equivariant puzzle piece such that $wt(p_{s})=y_{a}-y_{c}$ for some $a>c$. Define $\sigma_{r}wt(p_{s}):=y_{n+1-a}-y_{n+1-c}$ for $n+1-a<n+1-c$ and $\sigma_{r}wt(P):=\prod_{s=1}^{|P|}\sigma_{r}wt(p_{s})$.
Lemma 5.7.
$\widetilde{\mathscr{C}}_{i~{}j}^{\ell}=\sum_{P~{}:~{}\partial P=\Delta_{\sigma_{r}\lambda^{i}~{}\sigma_{r}\lambda^{j}}^{\sigma_{r}\lambda^{\ell}}}\sigma_{r}wt(P).$
Proof.
Apply $\sigma_{r}$ from left in (5.1) we get
$(\sigma_{r}\widetilde{S}_{\lambda^{i}})(\sigma_{r}\widetilde{S}_{\lambda^{j}})$ $=\sum_{\lambda^{\ell}}\sigma_{r}\widetilde{C}_{i~{}j}^{\ell}(\sigma_{r}\widetilde{S}_{\lambda^{\ell}})$
$(\widetilde{\mathbb{S}}_{\sigma_{r}\lambda^{i}})(\widetilde{\mathbb{S}}_{\sigma_{r}\lambda^{j}})$ $=\sum_{\lambda^{\ell}}\sigma_{r}\widetilde{C}_{i~{}j}^{\ell}(\widetilde{\mathbb{S}}_{\sigma_{r}\lambda^{\ell}})$
${\mathscr{\widetilde{C}}}_{\sigma_{r}\lambda^{i}~{}\sigma_{r}\lambda^{j}}^{\sigma_{r}\lambda^{\ell}}$ $=\sigma_{r}{\widetilde{C}}_{i~{}j}^{\ell}.$
Now the proof follows using Proposition 5.3. ∎
Next we use Lemma 5.7 to calculate the equivariant structure constant ${\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{\ell}$ for any arbitrary Plücker weight vector ${\bf{b}}$.
Recall (*$(\mathbb{C}^{*})^{n+1}$ action on $Pl(k,n)$ from (2.6) which is given by the map
$\rho\colon(\mathbb{C}^{*})^{n+1}\to(\mathbb{C}^{*})^{m+1}$
defined by $\rho(t_{1},t_{2},\dots,t_{n},t)=(t\cdot t_{\lambda^{0}},t\cdot t_{\lambda^{1}},\dots,t\cdot t_{\lambda^{m}}).$
Consider a weighted Grassmann orbifold $\mbox{Gr}_{\bf b}(k,n)$. Using Proposition 2.9, there exists $W=(w_{1},\dots,w_{n})\in\mathbb{Z}^{n}$ and $a\in\mathbb{Z}_{\geq 1}$ such that $b_{i}=w_{\lambda^{i}}$. Note that
$\rho(WD)=\{(t^{b_{0}},t^{b_{1}},\dots,t^{b_{m}})~{}|~{}t\in\mathbb{C}^{*}\},$
where $WD$ is the subgroup of (*$(\mathbb{C}^{*})^{n+1}$ defined in (2.7). Thus the action of the subgroup $\rho(WD)$ on $\rho((\mathbb{C}^{*})^{n+1})$ is same as the ${\bf b}$-action of $\mathbb{C}^{*}$. Let $T^{n+1}=(S^{1})^{n+1}$ and define $T_{\bf b}:=\frac{\rho(T^{n+1})}{{\bf b}\text{-action of }S^{1}}$. Consequently there is a $T_{\bf b}$ action on $\mbox{Gr}_{\bf{b}}(k,n)$. Note that $H^{*}(BT_{{\bf{b}}};\mathbb{Q})=\mathbb{Q}[x_{1},\dots,x_{n}]$, where $x_{1}=y_{1}-\frac{w_{1}}{a}z,\dots,x_{n}=y_{n}-\frac{w_{n}}{a}z$.
Using Vitoris Begle mapping theorem, the authors in [1, Proposition 3.1] proved that there is an isomorphism between the two equivariant cohomology ring with rational coefficients
$\pi^{*}:H_{T^{n}}^{*}(\mbox{Gr}(k,n);\mathbb{Q})\to H^{*}_{T^{n+1}}(Pl(k,n);\mathbb{Q})$
as rings over the polynomial rings $H^{*}(BT^{n};\mathbb{Q})$.
The advantage of taking the $T_{\bf{b}}$ action on $\mbox{Gr}_{\bf{b}}(k,n)$ is
$\pi^{*}_{{\bf{b}}}:H_{T_{\bf{b}}}^{*}(\mbox{Gr}_{\bf{b}}(k,n);\mathbb{Q})\to H^{*}_{T^{n+1}}(Pl(k,n);\mathbb{Q})$
is isomorphism as rings over the polynomial rings $H^{*}(BT_{\bf{b}};\mathbb{Q})$.
The authors in [1] computed the $T^{n+1}$-equivariant cohomology ring of $Pl(k,n)$ with rational coefficients. We denote $H^{*}(BT^{n+1};\mathbb{Q})=\mathbb{Q}[y_{1},\dots,y_{n},z]$.
Proposition 5.8.
[1, Theorem 1.1] The $T^{n+1}$-equivariant cohomology ring of $Pl(k,n)$ with rational coefficients is given by
$H_{T^{n+1}}^{*}(Pl(k,n);\mathbb{Q})=\Big{\{}\alpha\in\bigoplus_{i=0}^{m}\mathbb{Q}[y_{1},\dots,y_{n},z]/(Y_{\lambda^{i}}+z)~{}\big{|}~{}$ $(\alpha_{i}-\alpha_{j})\in(Y_{\lambda^{i}}+z,Y_{\lambda^{j}}+z)$
[…]
Note that there exists an isomorphism
$\phi^{*}:H^{*}(BT_{\bf{b}};\mathbb{Q})\to H^{*}(BT^{n};\mathbb{Q})$
defined by $\phi^{*}(x_{i})=y_{i}$. Observe that $\phi^{*}(f)=f|_{z=0}\in\mathbb{Q}[y_{1},\dots,y_{n}]$ for a polynomial $f\in\mathbb{Q}[x_{1},\dots,x_{n}]$.
Proposition 5.9.
The $T_{\bf{b}}$-equivariant cohomology of $\mbox{Gr}_{\bf{b}}(k,n)$ with rational coefficient is given by
$H_{T_{\bf{b}}}^{*}(\mbox{Gr}_{\bf{b}}(k,n);\mathbb{Q})=\Big{\{}\alpha\in\bigoplus_{j=0}^{m}\mathbb{Q}[x_{1},\dots,x_{n}]~{}\big{|}~{}(X_{\lambda^{i}}-d_{ij}X_{\lambda^{j}})~{}\mbox{divides}~{}(\alpha_{j}-\alpha_{i})\text{ if }\lambda^{i}\in R(\lambda^{j})\Big{\}}.$
Here $X_{\lambda}=\sum_{i=1}^{k}x_{\lambda_{i}}$ for every Schubert symbol $\lambda=(\lambda_{1},\dots,\lambda_{k})$.
For every $i\in\{0,1\dots,m\}$, define $p\widetilde{\mathbb{S}}_{\lambda^{i}}\in H^{*}_{T^{n+1}}(Pl(k,n);\mathbb{Q})$ by
(5.3) $p\widetilde{\mathbb{S}}_{\lambda^{i}}:=\pi^{*}(\widetilde{\mathbb{S}}_{\lambda^{i}}).$
Since $\pi^{*}$ is a ring isomorphism, $\{p\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{m}$ is a basis for $H^{*}_{T^{n+1}}(Pl(k,n);\mathbb{Q})$. Note that $f\in\mathbb{Q}[y_{1},\dots,y_{n},z]$ then $f|_{z=0}\in\mathbb{Q}[y_{1},\dots,y_{n}]$.
Lemma 5.10.
${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}=(\pi^{*}_{{\bf{b}}})^{-1}(p\widetilde{\mathbb{S}}_{\lambda^{i}}|_{z=0}).$
Proof.
$\widetilde{\mathbb{S}}_{\lambda^{i}}|_{\lambda^{i}}=\prod_{\lambda^{\ell}\in R(\lambda^{i})}(Y_{\lambda^{\ell}}-Y_{\lambda^{i}})$. Thus
$p\widetilde{\mathbb{S}}_{\lambda^{i}}|_{\lambda^{i}}=\prod_{\lambda^{\ell}\in R(\lambda^{i})}(Y_{\lambda^{\ell}}-Y_{\lambda^{i}})=\prod_{\lambda^{\ell}\in R(\lambda^{i})}(Y_{\lambda^{\ell}}+z)\in\mathbb{Q}[y_{1},y_{2},\dots,y_{n},z]/(Y_{\lambda^{i}}+z).$
Note that if $f\in\frac{\mathbb{Q}[y_{1},\dots,y_{n},z]}{(Y_{\lambda^{i}}+z)}$ then $f|_{z=0}\in\frac{\mathbb{Q}[y_{1},\dots,y_{n}]}{(Y_{\lambda^{i}})}$. Thus
$(p\widetilde{\mathbb{S}}_{\lambda^{i}}|_{\lambda^{i}})|_{z=0}=\prod_{\lambda^{\ell}\in R(\lambda^{i})}Y_{\lambda^{\ell}}\in\frac{\mathbb{Q}[y_{1},\dots,y_{n}]}{(Y_{\lambda^{i}})}.$
[…]
Since $(\pi_{b})^{*}$ is an isomorphism, $p\widetilde{\mathbb{S}}_{\lambda^{i}}|_{\lambda^{j}}\neq 0\implies\lambda^{j}\succeq\lambda^{i}$. Now the proof follows from the uniqueness of ${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}$. ∎
The set $\{{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{m}$ forms a basis for the equivariant cohomology ring $H_{T^{n}}^{*}(\mbox{Gr}_{\bf{b}}(k,n);\mathbb{Q})$ as well. To compute the equivariant structure constants with respect to the basis $\{{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{m}$, we utilize Lemma 5.7, along with the isomorphisms $\pi^{*}$ and $\pi_{\bf{b}}^{*}$, as well as Lemma 5.10. A priory, these constants are expected to be elements of $\mathbb{Q}[y_{1},y_{2},\dots,y_{n}]$. However, we proceed to establish that all the equivariant structure constants lie in $\mathbb{Z}[y_{1},y_{2},\dots,y_{n}]$ for a divisive weighted Grassmann orbifold.
Using Lemma 5.7 and (5.3) we have
(5.4) $p\widetilde{\mathbb{S}}_{\lambda^{i}}p\widetilde{\mathbb{S}}_{\lambda^{j}}=\sum_{\lambda^{q}\succeq\lambda^{i},\lambda^{j}}\sum_{P:\partial P=\Delta_{\sigma_{r}\lambda^{i}~{}\sigma_{r}\lambda^{j}}^{\sigma_{r}\lambda^{q}}}\sigma_{r}wt(P)p\widetilde{\mathbb{S}}_{\lambda^{q}},$
where
$\sigma_{r}wt(P)=\prod_{s=1}^{|P|}\sigma_{r}wt(p_{s})$
and $p_{s}$ ranges over the equivariant pieces on $P$ such that $\partial P=\Delta_{\sigma_{r}\lambda^{i}~{}\sigma_{r}\lambda^{j}}^{\sigma_{r}\lambda^{q}}$. Observe that $|P|=d(\lambda^{i})+d(\lambda^{j})-d(\lambda^{q})$ for a puzzle $P$ such that $\partial P=\Delta_{\sigma_{r}\lambda^{i}~{}\sigma_{r}\lambda^{j}}^{\sigma_{r}\lambda^{q}}$. Note that $\sigma_{r}wt(p_{s})=Y_{\lambda^{e}}-Y_{\lambda^{f}}$ for some $\lambda^{e}\prec\lambda^{f}$. We define
(5.5) ${\bf{b}}(p_{s}):=b_{e}-b_{f}.$
Proposition 5.11.
$p\widetilde{\mathbb{S}}_{\lambda^{1}}|_{\lambda^{i}}=(Y_{\lambda^{0}}+z)$ for $i=1,2,\dots,m$.
Proof.
Using [7, Lemma 5.6], we have
$\widetilde{\mathbb{S}}_{\lambda^{1}}|_{\lambda^{i}}=(Y_{\lambda^{0}}-Y_{\lambda^{i}}).$
Now, $\pi^{*}$ is a ring isomorphism and $p\widetilde{\mathbb{S}}_{\lambda^{1}}=\pi^{*}(\widetilde{\mathbb{S}}_{\lambda^{1}}).$ Thus
$p\widetilde{\mathbb{S}}_{\lambda^{1}}|_{\lambda^{i}}=Y_{\lambda^{0}}-Y_{\lambda^{i}}=Y_{\lambda^{0}}+z\in\frac{\mathbb{Q}[y_{1},\dots,y_{n},z]}{(Y_{\lambda^{i}}+z)}.$
Recall that $H^{*}_{T^{n+1}}(Pl(k,n);\mathbb{Q})$ is also a $\mathbb{Q}[x_{1},\dots,x_{n}]$ module and $(\pi_{\bf{b}})^{*}$ is a ring isomorphism over $\mathbb{Q}[x_{1},\dots,x_{n}]$ module. Next, we relate two different module structures on $H^{*}_{T^{n+1}}(Pl(k,n);\mathbb{Q})$. Note that $X_{\lambda^{i}}=Y_{\lambda^{i}}-\frac{b_{i}-a}{a}z$ for all $i=0,1,\dots,m$. Thus
$Y_{\lambda^{e}}-Y_{\lambda^{f}}=X_{\lambda^{e}}-X_{\lambda^{f}}+\frac{b_{e}-b_{f}}{a}z=X_{\lambda^{e}}-X_{\lambda^{f}}+\frac{b_{e}-b_{f}}{b_{0}}(Y_{\lambda^{0}}-X_{\lambda^{0}}+z).$
Using Proposition 5.11, define
$K:=X_{\lambda^{e}}-X_{\lambda^{f}}-\frac{b_{e}-b_{f}}{b_{0}}X_{\lambda^{0}}+\frac{b_{e}-b_{f}}{b_{0}}p\widetilde{\mathbb{S}}_{\lambda^{1}}\in H^{*}_{T^{n+1}}(Pl(k,n);\mathbb{Q}).$
Thus $K\cdot p\widetilde{\mathbb{S}}_{\lambda^{q}}\in H^{*}_{T^{n+1}}(Pl(k,n);\mathbb{Q})$ and
$\sigma_{r}wt(p_{s})\cdot p\widetilde{\mathbb{S}}_{\lambda^{q}}=K\cdot p\widetilde{\mathbb{S}}_{\lambda^{q}}\text{ for all }q=0,1,\dots,m.$
Then evaluating at $z=0$ on both sides we get the following relation in $H^{*}_{T^{n+1}}(Pl(k,n);\mathbb{Q})$.
$\sigma_{r}wt(p_{s})(p\widetilde{\mathbb{S}}_{\lambda^{q}}|_{z=0})={\bf{b}}wt(p_{s})(p\widetilde{\mathbb{S}}_{\lambda^{q}}|_{z=0})+\frac{{\bf{b}}(p_{s})}{b_{0}}(p\widetilde{\mathbb{S}}_{\lambda^{1}}|_{z=0})(p\widetilde{\mathbb{S}}_{\lambda^{q}}|_{z=0}),$
where ${\bf{b}}wt(p_{s})=(Y_{\lambda^{e}}-Y_{\lambda^{f}}-\frac{{\bf{b}}(p_{s})}{b_{0}}Y_{\lambda^{0}})$. Thus
(5.6) $\prod_{s=1}^{|P|}\sigma_{r}wt(p_{s})(p\widetilde{\mathbb{S}}_{\lambda^{q}}|_{z=0})$ $=\prod_{s=1}^{|P|}({\bf{b}}wt(p_{s})+\frac{{\bf{b}}(p_{s})}{b_{0}}(p\widetilde{\mathbb{S}}_{\lambda^{1}}|_{z=0}))(p\widetilde{\mathbb{S}}_{\lambda^{q}}|_{z=0})$
$=\sum_{s=0}^{|P|}a^{(P)}_{s}(p\widetilde{\mathbb{S}}_{\lambda^{1}}|_{z=0})^{s}(p\widetilde{\mathbb{S}}_{\lambda^{q}}|_{z=0}),$
where $a^{(P)}_{s}$ can be calculated using the following remark.
Remark 5.12.
If $\prod_{s=1}^{t}(a_{s}+b_{s}x)=\sum_{s=0}^{t}\alpha_{s}x^{s}$ then
$\alpha_{s}=\sum_{1\leq i_{1}<i_{2}\dots<i_{s}\leq t}b_{i_{1}}b_{i_{2}}\dots b_{i_{s}}\frac{a_{1}a_{2}\dots a_{t}}{a_{i_{1}}a_{i_{2}}\dots a_{i_{s}}}.$
Then using (5.4), (5.6) and Lemma 5.10, we have
${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{j}}$ $=\sum_{\lambda^{q}\succeq\lambda^{i},\lambda^{j}}\sum_{P:\partial P=\Delta_{\sigma_{r}\lambda^{i}~{}\sigma_{r}\lambda^{j}}^{\sigma_{r}\lambda^{q}}}\sum_{s=0}^{|P|}a^{(P)}_{s}({\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{1}})^{s}{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{q}}.$
We have the weighted Pieri rule for the multiplication with the equivariant Schubert class ${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{1}}$. For any two Schubert symbol $\lambda^{i}$ and $\lambda^{j}$ we denote $\lambda^{j}\to\lambda^{i}$ if $d(\lambda^{j})=d(\lambda^{i})+1$ and $\lambda^{i}\preceq\lambda^{j}$.
Proposition 5.13 (Equivariant Pieri rule).
[7, Proposition 5.7]
${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{1}}{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}=(Y_{\lambda^{0}}-\frac{b_{0}}{b_{i}}Y_{\lambda^{i}}){\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}+\sum_{\lambda^{j}:\lambda^{j}\to\lambda^{i}}\frac{b_{0}}{b_{i}}{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{j}}.$
For $s\geq 1$, let ${\bf{b}}\tilde{\mathscr{C}}_{q~{}g^{s}}^{\ell}\in\mathbb{Z}[y_{1},y_{2},\dots,y_{n}]$ which satisfies
(5.7) $({\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{1}})^{s}{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{q}}=\sum_{\lambda^{\ell}\succeq\lambda^{q}}{\bf{b}}\widetilde{\mathscr{C}}_{q~{}g^{s}}^{\ell}{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{\ell}}.$
Now evaluating the Proposition 5.13 $s$ times for $s\geq 1$, we can compute the equivariant structure constants ${\bf{b}}\tilde{\mathscr{C}}_{q~{}g^{s}}^{\ell}$ by the following:
${\bf{b}}\widetilde{\mathscr{C}}_{q~{}g^{s}}^{\ell}=\sum_{\begin{subarray}{c}\lambda^{\ell}=\lambda^{\ell_{0}}\to\lambda^{\ell_{1}}\to\\ \dots\to\lambda^{\ell_{r}}=\lambda^{q}\end{subarray}}\frac{(b_{0})^{r}}{b_{\ell_{1}}b_{\ell_{2}}\dots b_{\ell_{r}}}\sum_{J}(\prod_{t=0}^{r}(Y_{\lambda^{0}}-\frac{b_{0}}{b_{\ell_{t}}}Y_{\lambda^{\ell_{t}}})^{j_{t}}),$
where $J$ runs over all sequences $(j_{0},j_{1},\dots,j_{r})$ of non negative integers satisfying $j_{0}+j_{1}+\dots+j_{r}=s-r$. Note that $r=d(\lambda^{\ell})-d(\lambda^{q})$. Thus ${\bf{b}}\widetilde{\mathscr{C}}_{q~{}g^{s}}^{\ell}=0$ if $s<d(\lambda^{\ell})-d(\lambda^{q})$.
Theorem 5.14.
The equivariant structure constant ${\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{\ell}$ is given by
${\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{\ell}=\sum_{\lambda^{\ell}\succeq\lambda^{q}\succeq\lambda^{i},\lambda^{j}}\sum_{P:\partial P=\Delta_{\sigma_{r}\lambda^{i}~{}\sigma_{r}\lambda^{j}}^{\sigma_{r}\lambda^{q}}}\sum_{s=0}^{|P|}a^{(P)}_{s}{\bf{b}}\widetilde{\mathscr{C}}_{q~{}g^{s}}^{\ell}.$
Proof.
we have
${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{j}}$ $=\sum_{\lambda^{q}\succeq\lambda^{i},\lambda^{j}}\left(\sum_{P:\partial P=\Delta_{\sigma_{r}\lambda^{i}~{}\sigma_{r}\lambda^{j}}^{\sigma_{r}\lambda^{q}}}\sum_{s=0}^{|P|}a^{(P)}_{s}({\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{1}})^{s}\right){\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{q}}$
(5.8) $=\sum_{\lambda^{q}\succeq\lambda^{i},\lambda^{j}}\sum_{P:\partial P=\Delta_{\sigma_{r}\lambda^{i}~{}\sigma_{r}\lambda^{j}}^{\sigma_{r}\lambda^{q}}}\sum_{s=0}^{|P|}a^{(P)}_{s}\left(\sum_{\lambda^{\ell}\succeq\lambda^{q}}{\bf{b}}\widetilde{\mathscr{C}}_{q~{}g^{s}}^{\ell}{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{\ell}}\right)$
$=\sum_{\lambda^{\ell}}\left(\sum_{\lambda^{\ell}\succeq\lambda^{q}\succeq\lambda^{i},\lambda^{j}}\sum_{P:\partial P=\Delta_{\sigma_{r}\lambda^{i}~{}\sigma_{r}\lambda^{j}}^{\sigma_{r}\lambda^{q}}}\sum_{s=0}^{|P|}a^{(P)}_{s}{\bf{b}}\widetilde{\mathscr{C}}_{q~{}g^{s}}^{\ell}\right){\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{\ell}}$
Thus to calculate the equivariant structure constant ${\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{\ell}$ first we need to find and sum over all $\lambda^{q}$ which satisfies ${\lambda^{\ell}\succeq\lambda^{q}\succeq\lambda^{i},\lambda^{j}}$. For each such $\lambda^{q}$ we need to sum over all equivariant puzzle $P$ such that $\partial P=\Delta_{\sigma_{r}\lambda^{i}~{}\sigma_{r}\lambda^{j}}^{\sigma_{r}\lambda^{q}}$. The number of equivariant puzzle pieces of $P$ is $|P|=d(\lambda^{i})+d(\lambda^{j})-d(\lambda^{q})$.
Now we need to prove that ${\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{\ell}\in\mathbb{Z}[y_{1},y_{2},\dots,y_{n}]$ for all $i,j$ and $\ell$. The argument follows by using ${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}|_{\lambda^{i}}$ is a primitive polynomial in $\mathbb{Z}[y_{1},y_{2},\dots,y_{n}]$ and the upper triangularity of the equivariant Schubert classes ${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}$. We recall a polynomial in $\mathbb{Z}[y_{1},y_{2},\dots,y_{n}]$ is said to be primitive if the gcd of all coefficients of the monomials is one.
Theorem 5.15.
Let ${\rm Gr}_{\bf{b}}(k,n)$ be a divisive weighted Grassmann orbifold. Then the equivariant structure constants ${\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{\ell}$ are in $\mathbb{Z}[y_{1},y_{2},\dots,y_{n}]$ for all $i,j,\ell\in\{0,1,\dots,m\}$.
Proof.
Let $\lambda^{q}$ be a minimal Schubert symbol such that $\lambda^{i}\preceq\lambda^{q}$ and $\lambda^{j}\preceq\lambda^{q}$ both holds Here $\lambda^{q}$ is minimal means there is no other Schubert symbol $\lambda^{p}$ such that $\lambda^{p}\prec\lambda^{q}$ and $\lambda^{i}\preceq\lambda^{p}$, $\lambda^{j}\preceq\lambda^{p}$ both hold. Therefore, $({\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{j}})|_{\lambda^{\ell}}\neq 0$ implies $\lambda^{\ell}\succeq\lambda^{q}$ and consequently ${\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{\ell}\neq 0$ implies $\lambda^{\ell}\succeq\lambda^{q}$. Note that if $\lambda^{\ell}\succ\lambda^{q}$ then ${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{\ell}}|_{\lambda^{q}}=0$. Considering the $\lambda^{q}$-th component in the both the side of (5.2) we have
(5.9) $({\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{j}})|_{\lambda^{q}}={\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{q}{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{q}}|_{\lambda^{q}}.$
Now $({\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{j}})|_{\lambda^{q}}\in\mathbb{Z}[y_{1},y_{2},\dots,y_{n}]$ and
${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{q}}|_{\lambda^{q}}=\prod_{\lambda^{\ell}\in R(\lambda^{q})}(Y_{\lambda^{\ell}}-\dfrac{b_{\ell}}{b_{q}}Y_{\lambda^{q}})\in\mathbb{Z}[y_{1},y_{2},\dots,y_{n}].$
Using Theorem 5.14, ${\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{q}\in\mathbb{Q}[y_{1},y_{2},\dots,y_{n}]$. If ${\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{q}\notin\mathbb{Z}[y_{1},y_{2},\dots,y_{n}]$, let $d$ be the product of the denominators of the coefficients of the monomials in the expression of ${\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{q}$. Then $\overline{\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{q}:=d{\cdot}{\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{q}\in\mathbb{Z}[y_{1},\dots,y_{n}]$.
Now multiplying both sides of (5.9) by $d$ we get
(5.10) $d{\cdot}({\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{j}})|_{\lambda^{q}}=\overline{\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{q}\prod_{\lambda^{\ell}\in R(\lambda^{q})}(Y_{\lambda^{\ell}}-\dfrac{b_{\ell}}{b_{q}}Y_{\lambda^{q}})$
Let $p$ be a prime divisor of $d$. Then reducing (5.10) modulo $p$, we obtain the following:
$0=\mathscr{P}_{1}\mathscr{P}_{2}~{}~{}~{}~{}(\text{mod}~{}p),$
where $\mathscr{P}_{1}=\overline{\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{q}~{}~{}~{}~{}(\text{mod}~{}p)$ and $\mathscr{P}_{2}=\prod_{\lambda^{\ell}\in R(\lambda^{q})}(Y_{\lambda^{\ell}}-\dfrac{b_{\ell}}{b_{q}}Y_{\lambda^{q}})~{}~{}~{}~{}(\text{mod}~{}p)$.
Note that $\prod_{\lambda^{\ell}\in R(\lambda^{q})}(Y_{\lambda^{\ell}}-\dfrac{b_{\ell}}{b_{q}}Y_{\lambda^{q}})$ is a primitive polynomial. Thus $\mathscr{P}_{2}\neq 0~{}~{}~{}~{}(\text{mod}~{}p)$. Also $({\mathbb{Z}}/{p\mathbb{Z}})[y_{1},\dots,y_{n}]$ is an integral domain. Thus $\mathscr{P}_{1}=0~{}~{}~{}~{}(\text{mod}~{}p)$. Thus $\frac{1}{p}\overline{\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{q}$ is in $\mathbb{Z}[y_{1},\dots,y_{n}]$. In other words, we can cancel a factor of $p$ in $d$ in (5.10), and still have an equation in $\mathbb{Z}[y_{1},\dots,y_{n}]$. Now using induction on the prime factors of $d$ we get
${\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{q}=\frac{1}{d}\overline{\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{q}\in\mathbb{Z}[y_{1},\dots,y_{n}].$
Next, consider the minimal Schubert symbol $\lambda^{r}$ such that
$({\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{j}}-{\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{q}{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{q}})|_{\lambda^{r}}\neq 0.$
Then we follow the same argument to conclude
${\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{r}\in\mathbb{Z}[y_{1},y_{2},\dots,y_{n}].$
Now we can complete the proof inductively. ∎
Next, we prove the positivity condition of the equivariant structure constants ${\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{\ell}$. In [9], the author shows that $\widetilde{{C}}_{i~{}j}^{\ell}$ is a polynomial of $y_{q+1}-y_{q}$ with non negative coefficients for $q=1,2,\dots,n-1$.
Theorem 5.16.
The equivariant structure constant ${\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{\ell}$ can be written as a polynomial of $y_{q}-y_{q+1}-\frac{b_{u}-b_{v}}{b_{0}}Y_{\lambda^{0}}$ with non negative coefficients for $q=1,2,\dots,{n-1}$, where $\lambda^{u}=(q+1,q)\lambda^{v}$ for some $u<v$.
Proof.
Recall that ${\bf{b}}wt(p_{s})=Y_{\lambda^{e}}-Y_{\lambda^{f}}-\frac{b_{e}-b_{f}}{b_{0}}Y_{\lambda^{0}}$. Let $Y_{\lambda^{e}}-Y_{\lambda^{f}}=y_{r}-y_{s}$ for some $1\leq r<s\leq n$. Then ${\bf{b}}wt(p_{s})$ can be written as the linear sum of the polynomials $(y_{q}-y_{q+1}-\frac{b_{u}-b_{v}}{b_{0}}Y_{\lambda^{0}})$, for $q=r,r+1,\dots,s$. Also $b_{u}-b_{v}>0$ as ${\bf{b}}$ is a divisive Plücker weight vector. Thus, the term $a^{(P)}_{s}$ in the statement of the Theorem 5.14 satisfies the claim. Now we need to prove the same for ${\bf{b}}\widetilde{\mathscr{C}}_{q~{}g^{s}}^{\ell}$. Note that if $\lambda^{\ell^{\prime}}=(c+1,c)\lambda^{\ell}$ for some $(c+1,c)\in\text{rev}(\lambda^{\ell})$ then $\lambda^{\ell^{\prime}}<\lambda^{\ell}$ and
$(Y_{\lambda^{0}}-\frac{b_{0}}{b_{\ell}}Y_{\lambda^{\ell}})=\frac{b_{\ell^{\prime}}}{b_{\ell}}(Y_{\lambda^{0}}-\frac{b_{0}}{b_{\ell^{\prime}}}Y_{\lambda^{\ell^{\prime}}})+\frac{b_{0}}{b_{\ell}}(y_{c}-y_{c+1}-\frac{b_{\ell^{\prime}}-b_{\ell}}{b_{0}}Y_{\lambda^{0}}).$
Also $Y_{\lambda^{0}}-\frac{b_{0}}{b_{1}}Y_{\lambda^{1}}=\frac{b_{0}}{b_{1}}(Y_{\lambda^{0}}-Y_{\lambda^{1}}-\frac{b_{1}-b_{0}}{b_{0}}Y_{\lambda^{0}})$. This completes the proof. ∎
6. Structure constants for the integral cohomology ring of divisive weighted Grassmann orbifolds
In this section, we compute the integral cohomology rings of divisive weighted Grassmann orbifolds. First, we describe the integral cohomology groups of divisive weighted Grassmann orbifolds using Theorem 4.12. Then, we consider a basis for the cohomology groups and compute the cup product concerning that basis.
Proposition 6.1.
Let $\mbox{Gr}_{\bf{b}}(k,n)$ be a divisive weighted Grassmann orbifold. Then
$H^{j}(\mbox{Gr}_{b}(k,n);\mathbb{Z})=\begin{cases}\bigoplus_{i:2d(\lambda^{i})=j}\mathbb{Z}&\quad\text{if}~{}j\text{ is even }\\ 0&\quad\text{otherwise}.\end{cases}$
For every Schubert symbol $\lambda^{i}$, we can define the weighted Schubert class
${\bf{b}}{\mathbb{S}}_{\lambda^{i}}:=(\xi)^{*}({\bf{b}}{\widetilde{\mathbb{S}}}_{\lambda^{i}}),$
where $\xi$ is the inclusion map
$\xi\colon\mbox{Gr}_{\bf{b}}(k,n)\to ET^{n}\times_{T^{n}}\mbox{Gr}_{\bf{b}}(k,n).$
Geometrically, ${\bf b}\widetilde{\mathbb{S}}_{\lambda^{i}}$ is the equivariant cohomology class in the integral equivariant cohomology ring of the divisive weighted Grassmann orbifold corresponding to the closure of the cell $\pi_{\bf b}(\sigma_{r}\widetilde{E}(\sigma_{r}\lambda^{i}))$. Thus $(\xi)^{*}({\bf b}\widetilde{\mathbb{S}}_{\lambda^{i}})$ represents the Poincare dual of the cell $\pi_{\bf b}(\sigma_{r}\widetilde{E}(\sigma_{r}\lambda^{i}))$ in ordinary cohomology. Note that dimension of $\pi_{\bf b}(\sigma_{r}\widetilde{E}(\sigma_{r}\lambda^{i}))$ is $2(k(n-k)-d(\lambda^{i}))$. Therefore, $\{{\bf{b}}{\mathbb{S}}_{\lambda^{i}}:2d(\lambda^{i})=j\}$ forms a basis for the ordinary cohomology group $H^{j}(\mbox{Gr}_{\bf{b}}(k,n);\mathbb{Z})$.
Let $\lambda^{i}$, $\lambda^{j}$ and $\lambda^{\ell}$ be three Schubert symbols such that $d(\lambda^{i})+d(\lambda^{j})=d(\lambda^{\ell})$. Then we have the following product formulae in the ordinary cohomology ring of divisive weighted Grassmann orbifold.
(6.1) ${\bf{b}}{\mathbb{S}}_{\lambda^{i}}{\bf{b}}{\mathbb{S}}_{\lambda^{j}}=\sum_{\lambda^{\ell}~{}:~{}d(\lambda^{\ell})=d(\lambda^{i})+d(\lambda^{j})}{\bf{b}}{\mathscr{C}}_{i~{}j}^{\ell}{\bf{b}}{\mathbb{S}}_{\lambda^{\ell}}.$
We make use of Theorem 5.14 to deduce a formulae for the structure constant ${\bf{b}}{\mathscr{C}}_{i~{}j}^{\ell}$. Note that the equivariant structure constant ${\bf{b}}{\widetilde{\mathscr{C}}}_{i~{}j}^{\ell}$ coincides with ${\bf{b}}{\mathscr{C}}_{i~{}j}^{\ell}$ if $d(\lambda^{i})+d(\lambda^{j})=d(\lambda^{\ell})$. We recall for any two Schubert symbol $\lambda^{i}$ and $\lambda^{j}$ we denote $\lambda^{j}\to\lambda^{i}$ if $d(\lambda^{j})=d(\lambda^{i})+1$ and $\lambda^{i}\preceq\lambda^{j}$. We have the following formulae.
Proposition 6.2 (weighted Pieri rule).
${\bf{b}}{\mathbb{S}}_{\lambda^{1}}{\bf{b}}{\mathbb{S}}_{\lambda^{i}}=\sum_{\lambda^{j}:\lambda^{j}\to\lambda^{i}}\frac{b_{0}}{b_{i}}{\bf{b}}{\mathbb{S}}_{\lambda^{j}}$
for all $i=1,2,\dots,m$.
Proof.
Proof follows from the proof of Proposition 5.13. ∎
Using Lemma 5.7 it follows $\widetilde{\mathscr{C}}_{i~{}j}^{q}$ is a polynomial of $Y_{\lambda^{e}}-Y_{\lambda^{f}}$ for some $\lambda^{e}\prec\lambda^{f}$. Define
$\widetilde{\mathscr{C}}_{i~{}j}^{q}({\bf{b}}):=\sum_{P~{}:~{}\partial P=\Delta_{\sigma_{r}\lambda^{i}~{}\sigma_{r}\lambda^{j}}^{\sigma_{r}\lambda^{q}}}\prod_{s=1}^{|P|}{\bf{b}}(p_{s}),$
where ${\bf{b}}(p_{s})$ is defined in (5.5) for every equivariant puzzle pieces $p_{s}$ of $P$ for $s=1,2,\dots,|P|$. For every Schubert symbol $\lambda^{q}$ such that $\lambda^{\ell}\succ\lambda^{q}\succeq\lambda^{i},\lambda^{j}$ and for every $\lambda^{\ell}=\lambda^{\ell_{0}}\rightarrow\lambda^{\ell_{1}}\rightarrow\dots\rightarrow\lambda^{\ell_{d}}=\lambda^{q}$ we can calculate
(6.2) $\mathscr{D}_{i~{}j}^{\ell~{}q}=\sum_{\begin{subarray}{c}\lambda^{\ell}=\lambda^{\ell_{0}}\rightarrow\lambda^{\ell_{1}}\rightarrow\\ \dots\rightarrow\lambda^{\ell_{d}}=\lambda^{q}\end{subarray}}\frac{\widetilde{\mathscr{C}}_{i~{}j}^{q}({\bf{b}})}{b_{\ell_{1}}b_{\ell_{2}}\dots b_{\ell_{d}}}.$
Then we have the following formulae for the structure constant ${\bf{b}}\mathscr{C}_{i~{}j}^{\ell}$ for the ordinary cohomology ring of divisive weighted Grassmann orbifold with integer coefficients.
Theorem 6.3.
The structure constant ${\bf{b}}\mathscr{C}_{i~{}j}^{\ell}$ is given by
${\bf{b}}\mathscr{C}_{i~{}j}^{\ell}=\mathscr{C}_{i~{}j}^{\ell}+\sum_{\lambda^{q}~{}:~{}\lambda^{\ell}\succ\lambda^{q}\succeq\lambda^{i},\lambda^{j}}\mathscr{D}_{i~{}j}^{\ell~{}q}.$
Proof.
We recall the formulae for equivariant structure constant ${\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{\ell}$ from Theorem 5.14. The maximum value of $d(\lambda^{\ell})$ which satisfy (5.7) is $d(\lambda^{q})+s$. Thus the maximum value of $d(\lambda^{\ell})$ which satisfy (5.2) is $d(\lambda^{q})+|P|$. Now $|P|=d(\lambda^{i})+d(\lambda^{j})-d(\lambda^{q})$. Thus $d(\lambda^{\ell})\leq d(\lambda^{i})+d(\lambda^{j})$. The equivariant structure constant ${\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{\ell}$ becomes the structure constant ${\bf{b}}\mathscr{C}_{i~{}j}^{\ell}$ for the ordinary cohomology if $d(\lambda^{i})+d(\lambda^{j})=d(\lambda^{\ell})$. Thus we need to consider $s=|P|=d(\lambda^{i})+d(\lambda^{j})-d(\lambda^{q})=d(\lambda^{\ell})-d(\lambda^{q}).$ Let $d:=d(\lambda^{\ell})-d(\lambda^{q})$.
${\bf{b}}{\mathscr{C}}_{i~{}j}^{\ell}$ $=\sum_{\lambda^{\ell}\succeq\lambda^{q}\succeq\lambda^{i},\lambda^{j}}\sum_{P:\partial P=\Delta_{\sigma_{r}\lambda^{i}~{}\sigma_{r}\lambda^{j}}^{\sigma_{r}\lambda^{q}}}a^{(P)}_{d}{\bf{b}}\widetilde{\mathscr{C}}_{q~{}g^{d}}^{\ell}$
$=\sum_{\lambda^{\ell}\succeq\lambda^{q}\succeq\lambda^{i},\lambda^{j}}\sum_{P:\partial P=\Delta_{\sigma_{r}\lambda^{i}~{}\sigma_{r}\lambda^{j}}^{\sigma_{r}\lambda^{q}}}\sum_{\begin{subarray}{c}\lambda^{\ell}=\lambda^{\ell_{0}}\to\lambda^{\ell_{1}}\to\\ \dots\to\lambda^{\ell_{d}}=\lambda^{q}\end{subarray}}\frac{(b_{0})^{d}}{b_{\ell_{1}}b_{\ell_{2}}\dots b_{\ell_{d}}}\prod_{s=1}^{d}(\frac{{\bf{b}}(p_{s})}{b_{0}})$
$=\sum_{\lambda^{\ell}\succeq\lambda^{q}\succeq\lambda^{i},\lambda^{j}}\sum_{\begin{subarray}{c}\lambda^{\ell}=\lambda^{\ell_{0}}\to\lambda^{\ell_{1}}\to\\ \dots\to\lambda^{\ell_{d}}=\lambda^{q}\end{subarray}}\frac{1}{b_{\ell_{1}}b_{\ell_{2}}\dots b_{\ell_{d}}}\sum_{P:\partial P=\Delta_{\sigma_{r}\lambda^{i}~{}\sigma_{r}\lambda^{j}}^{\sigma_{r}\lambda^{q}}}\prod_{s=1}^{d}{\bf{b}}(p_{s})$
$=\sum_{\lambda^{\ell}\succeq\lambda^{q}\succeq\lambda^{i},\lambda^{j}}\sum_{\begin{subarray}{c}\lambda^{\ell}=\lambda^{\ell_{0}}\to\lambda^{\ell_{1}}\to\\ \dots\to\lambda^{\ell_{d}}=\lambda^{q}\end{subarray}}\frac{\widetilde{\mathscr{C}}_{i~{}j}^{q}({\bf{b}})}{b_{\ell_{1}}b_{\ell_{2}}\dots b_{\ell_{d}}}$
$=\mathscr{C}_{i~{}j}^{\ell}+\sum_{\lambda^{q}~{}:~{}\lambda^{\ell}\succ\lambda^{q}\succeq\lambda^{i},\lambda^{j}}\mathscr{D}_{i~{}j}^{\ell~{}q}$
Using the above theorem, one can compute the integral cohomology ring of a divisive weighted Grassmann orbifold explicitly.
Remark 6.4.
For a divisive $\mbox{Gr}_{\bf b}(k,n)$ the structure constants ${\bf{b}}\mathscr{C}_{i~{}j}^{\ell}\in\mathbb{Z}$ and all are non-negative follows from Theorem 5.15 and Theorem 5.16 respectively.
In the following example, we compute the integral cohomology ring of a divisive weighted Grassmann orbifold $\mbox{Gr}_{\bf{b}}(2,4)$.
Example 6.5.
Consider a Plücker weight vector ${\bf{b}}:=(\alpha\beta,\alpha\beta,\alpha\beta,\alpha,\alpha,\alpha)$, where $\alpha$ and $\beta$ are two positive integers. Then ${\bf{b}}$ is a divisive Plücker weight vector and the corresponding weighted Grassmann orbifold $\mbox{Gr}_{\bf{b}}(2,4)$ is a divisive weighted Grassmann orbifold. In this case, we have $6$ Schubert symbols given by
$\lambda^{0}=(1,2)<\lambda^{1}=(1,3)<\lambda^{2}=(1,4)<\lambda^{3}=(2,3)<\lambda^{4}=(2,4)<\lambda^{5}=(3,4).$
Corresponding to every Schubert symbol we have $6$ basis element $\{{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{5}$. From Proposition 6.2 it follows that
${\bf{b}}{C}_{1~{}i}^{j}=\frac{b_{0}}{b_{i}}\text{ if }d(\lambda^{j})=d(\lambda^{i})+1\text{ and }\lambda^{j}\succ\lambda^{i}.$
Now we calculate the remaining structure constants. Note that from [14] we have
$\widetilde{{S}}_{\lambda^{3}}\widetilde{{S}}_{\lambda^{3}}$ $=(y_{4}-y_{2})(y_{4}-y_{3})\widetilde{{S}}_{\lambda^{3}}+(y_{4}-y_{3})\widetilde{{S}}_{\lambda^{1}}+\widetilde{{S}}_{\lambda^{0}}$
$=(Y_{\lambda^{5}}-Y_{\lambda^{3}})(Y_{\lambda^{2}}-Y_{\lambda^{1}})\widetilde{{S}}_{\lambda^{3}}+(Y_{\lambda^{2}}-Y_{\lambda^{1}})\widetilde{{S}}_{\lambda^{1}}+\widetilde{{S}}_{\lambda^{0}}.$
Applying $\sigma_{r}$ both side we get
$\widetilde{\mathbb{S}}_{\lambda^{3}}\widetilde{\mathbb{S}}_{\lambda^{3}}=(Y_{\lambda^{0}}-Y_{\lambda^{3}})(Y_{\lambda^{2}}-Y_{\lambda^{4}})\widetilde{\mathbb{S}}_{\lambda^{3}}+(Y_{\lambda^{2}}-Y_{\lambda^{4}})\widetilde{\mathbb{S}}_{\lambda^{4}}+\widetilde{\mathbb{S}}_{\lambda^{5}}.$
Then
${\bf{b}}\mathscr{C}_{3~{}3}^{5}$ $=\mathscr{C}_{3~{}3}^{5}+\mathscr{D}_{3~{}3}^{5~{}4}+\mathscr{D}_{3~{}3}^{5~{}3}$
$=1+\frac{b_{2}-b_{4}}{b_{4}}+\frac{(b_{0}-b_{3})(b_{2}-b_{4})}{b_{4}b_{3}}$
$=1+\frac{b_{2}-b_{4}}{b_{4}}(1+\frac{b_{0}-b_{3}}{b_{3}})$
$=1+\frac{(b_{2}-b_{4})b_{0}}{b_{4}b_{3}}=1+\beta(\beta-1).$
Similarly one can calculate
${\bf{b}}\mathscr{C}_{2~{}2}^{5}=1+\frac{b_{0}(b_{1}-b_{2})}{b_{2}b_{4}}=1+\beta(\beta-1).$
From [14]
$\widetilde{{S}}_{\lambda^{3}}\widetilde{{S}}_{\lambda^{2}}=(y_{4}-y_{1})\widetilde{{S}}_{\lambda^{1}}=(Y_{\lambda^{5}}-Y_{\lambda^{1}})\widetilde{{S}}_{\lambda^{1}}.$
Applying $\sigma_{r}$ both side we get
$\widetilde{\mathbb{S}}_{\lambda^{3}}\widetilde{\mathbb{S}}_{\lambda^{2}}=(Y_{\lambda^{0}}-Y_{\lambda^{4}})\widetilde{\mathbb{S}}_{\lambda^{4}}.$
Then
${\bf{b}}\mathscr{C}_{2~{}3}^{5}$ $=\mathscr{C}_{2~{}3}^{5}+\mathscr{D}_{2~{}3}^{5~{}4}$
$=0+\frac{b_{0}-b_{4}}{b_{4}}=\beta-1.$
All the remaining structure constants in the integral cohomology ring of the above divisive $\mbox{Gr}_{\bf{b}}(2,4)$ are zero. This gives the integral cohomology ring of the given divisive weighted Grassmann orbifold explicitly.
Acknowledgement. The author would like to express my special thanks of gratitude to Soumen Sarkar for his guidance and support. We have many valuable discussions. The author thanks Hiraku Abe and Tommo Matsumura for helpful discussions. The author would like to thank IIT Madras for the Ph.D. fellowship.